### v2

# finantial support acknowledgement

Abstract
Federated Learning, designed to address privacy concerns in learning models, introduces a new distributed paradigm that safeguards data privacy but differentiates the attack surface due to the server’s inaccessibility to local datasets and the change in protection objective–parameters’ integrity. Existing approaches, including robust aggregation algorithms, fail to effectively filter out malicious clients, especially those with non-Independently and Identically Distributed data. Furthermore, these approaches often tackle non-IID data and poisoning attacks separately. To address both challenges simultaneously, we present FedCC, a simple yet novel algorithm. It leverages the Centered Kernel Alignment similarity of Penultimate Layer Representations for clustering, allowing it to identify and filter out malicious clients by selectively averaging chosen parameters, even in non-IID data settings. Our extensive experiments demonstrate the effectiveness of FedCC in mitigating untargeted model poisoning and backdoor attacks. FedCC reduces the attack confidence to a consistent zero compared to existing outlier detection-based and first-order statistics-based methods. Specifically, it significantly minimizes the average degradation of global performance by 65.5%. We believe that this new perspective of assessing learning models makes it a valuable contribution to the field of FL model security and privacy. The code will be made available upon paper acceptance.
Index Terms:
Backdoor Attack, Defense, Federated Learning, Model Poisoning Attack, Non-IID, Robust Aggregation
[…]
=-21pt
I Introduction
Federated Learning (FL) [17] has emerged as a promising privacy-preserving model training approach in response to growing concerns about privacy breaches and data leakage caused by centralized learning. While Machine Learning (ML) and Deep Learning (DL) algorithms are popular for their capability, such as personalized recommendations and accurate forecasts, their centralized nature risks data privacy. FL allows data to remain on client devices but exchanges their local model parameters only, avoiding the need for central data collection.
FedAvg [17] is the first algorithm run on the server to aggregate clients’ model parameters. It involves a server initializing and advertising a global model to all clients, who then train the model with their private data and upload the computed parameters to the server. The server averages these parameters to update the global model. These steps are repeated until a stopping criterion is met. FL is especially suitable for data-sensitive environments since raw data remains on clients, reducing the risk of data leakage on the server or in-between communication.
However, due to its distributed nature, FL is susceptible to model poisoning attacks [12]. The server cannot directly examine local datasets’ data quality or model parameter integrity, allowing compromised clients or attackers to manipulate local models. The attacks can degrade global model performance indiscriminately (untargeted attacks) [8, 2, 18] or cause incorrect predictions on specific inputs (targeted attacks) [21, 23, 3, 25]. Backdoor attacks, a type of targeted attack, are stealthier by maintaining overall performance while misclassifying specific inputs [1].
Some preliminary defenses propose robust aggregation methods, but they may compromise privacy by sharing raw data or exposing data distribution to unreliable parties [6, 22]. Existing defenses often target IID data settings, which are rare in practice. Non-IID data, with variations in class distribution, features, or labels across clients [12], can gradually degrade model performance and complicate the detection of malicious clients [28, 27], as the impact of attacks increases with the degree of non-IID [16, 18].
In this study, we present FedCC, a novel defense mechanism for FL that mitigates untargeted model poisoning and targeted backdoor attacks, even with non-IID data. FedCC leverages the Centered Kernel Alignment (CKA) of Penultimate Layer Representations (PLRs) to distinguish between malicious and benign clients while maintaining data privacy by avoiding sharing raw data or additional information between the server and clients. PLRs contain highly distinguishable features, maximizing the differences in CKA scores between benign and malicious clients, even with non-IID data and no server knowledge of local data.
Empirical experiments show that FedCC outperforms existing defenses that rely on first-order statistics. It significantly improves global accuracy against untargeted attacks (65.5% compared to similar approaches) and reduces attack confidence in targeted backdoor attacks to almost zero while preserving main task accuracy. These results highlight FedCC’s effectiveness and robustness in mitigating attacks in FL settings, particularly in non-IID data settings.
Our contributions are summarized as follows:
  • We propose FedCC, a simple yet novel detection and aggregation method for FL, to defend against model poisoning attacks (untargeted model poisoning and targeted backdoor attacks) by comparing the CKA of PLRs of local and global models and averaging the selected local model parameters.
  • We justify that PLR is the most distinguishable and indicative feature to classify malicious and benign clients by measuring the distance between two clusters of clients.
  • We empirically show that CKA is the most sensitive and accurate metric to measure similarities between differently trained models, especially when data are non-IID, and the server does not have any knowledge about clients or any centralized data.
  • We empirically demonstrated the effectiveness of FedCC for defending against model poisoning attacks. Our experimental results show that comparing the CKA of PLRs is more reliable in filtering malicious clients out than using first-order statistics. FedCC outperforms existing defenses even when data are non-IID.
II Backgrounds and Related Works
II-A Poisoning Attacks in FL
FL is vulnerable to poisoning attacks due to its distributed nature and the server’s inability to examine dataset quality and model parameter integrity directly. Model poisoning is particularly destructive among various poisoning attacks, as adversaries stealthily manipulate local model parameters to maximize overall performance damage [3, 5]. Thus, this paper focuses on model poisoning attacks.
Model poisoning attacks allow client-side adversaries to alter local model parameters before sending them to the server. To avoid the global model from divergence, attackers optimize the compromised local model for both training loss and an adversarial objective, enhancing its stealthiness [5, 8]. They can also adjust model hyperparameters, such as the learning rate, local epochs, batch size, and optimization objective, to manipulate training rules before and during local training.
Based on the attacker’s goals, model poisoning attacks can be classified as untargeted or targeted. Untargeted attacks aim to degrade overall global test accuracy by manipulating compromised local models. In contrast, targeted attacks aim to mislead the model into misclassifying specific inputs, leaving other classes unaffected. This paper introduces a backdoor ability into targeted model poisoning attacks, where inputs belonging to a chosen class are intentionally misclassified to a chosen label, whereas the performance of other classes remains unaffected. Its added stealthiness makes the attack more challenging to detect [23, 21].
II-B Robust Aggregation in FL
Several Byzantine-robust aggregations based on summary statistics or anomaly detection have been proposed to mitigate model poisoning and backdoor attacks in FL. For example, Krum [4] selected the update with the smallest Euclidean distance to the remaining updates as the new global model. Still, it ignored outliers and assumed IID data only [2]. Median [26] selects the coordinate-wise median of updates, also ignoring outliers. Trimmed Mean [26] excludes extreme values before averaging, and Bulyan [11] combines Krum and Trimmed Mean for selection. These methods rely on knowing the number of attackers, which is often unavailable in practice.
Inspired by the fact that Sybil’s crafted model parameter’s directions are more similar to each other than to benign ones, Foolsgold [9] identified Sybils by measuring cosine similarity between local model gradients, marking the most similar ones as Sybils. However, it struggled with a single malicious client when there were no other malicious clients for comparison. Additionally, this approach may overfit non-IID data, ironically leading to suboptimal performance with IID data. FLTrust [6] used trust scores based on ReLU-clipped cosine similarities to weigh local clients’ parameters differently but risked privacy by requiring clients to share raw data with the server. Furthermore, precise weight scaling is paramount to ensure model convergence.
Previous works often used summary statistics (mean or median), direction (cosine similarity), or distance (Euclidean) of weight vectors, but they struggled to effectively distinguish between benign clients with non-IID data and malicious clients, often misclassifying benign clients as malicious. The lack of performance might be partially attributed to the inadequacy of Euclidean geometry as a metric to compare neural network representations. In contrast, our proposed approach utilizes kernel-based metric (Kernel CKA) to measure the similarity between global and local models, reliably identifying compromised model parameters among the mixture of clients and enhancing defense against model poisoning attacks in FL.
II-C Penultimate Layer Representation in FL
Figure 1: CNN architecture and Penultimate Layer
The penultimate layer is the second-to-last layer (before the softmax layer) of a neural network model, as illustrated in Figure 1. Wang et al. [22] discovered that PLRs are effective in distinguishing malicious models from benign ones; within benign clients, benign PLRs follow the same distribution, whereas malicious PLRs do not, across multiple datasets and neural network architectures. Specifically, they demonstrated that distances among benign PLRs are smaller than those of benign and malicious PLRs. Their proposed method, FLARE, assigned a trust score to each client based on pairwise PLR discrepancies defined by Maximum Mean Discrepancy (MMD) among all model updates, allocating lower values to those farther from the benign distribution. The model updates are then scaled and averaged, weighted by each client’s trust score. However, FLARE continuously redirects the global model using server-owned raw data, increasing the risk of data leakage and jeopardizing data privacy.
II-D Centered Kernel Alignment
Centered Kernel Alignment (CKA) [13] is a highly accurate similarity metric that measures how similar two differently initialized or trained neural networks are, producing a value between 0 (not similar) and 1 (identical). CKA can be calculated by:
$CKA(X,Y)$ $=\frac{HSIC(X,Y)}{\sqrt{HSIC(X,X)HSIC(Y,Y)}}$
$=\frac{{\lVert K(X^{T})K(Y)\rVert}^{2}}{{\lVert K(X^{T})K(X)\rVert}{\lVert K(Y^{T})K(Y)\rVert}}$
where $HSIC$ is Hilbert-Schmidt Independence Criterion [10] (i.e., a not normalized version of CKA) and $K$ is the RBF kernel. It constructs similarity kernel matrices in the weight space and compares these matrices to characterize the representation space, allowing comparisons between layers with varying widths and different initialization schemes.
CKA includes several important properties for similarity measures of neural networks. It is non-invariant to invertible linear transformations, meaning similarity scores change if such transformations are applied. This property is vital; the gradient descent algorithm is not invariant to these transformations because similarity metrics invariant to linear transformations are inaccurate on models trained with gradient descent. It is also invariant to orthogonal transformations and isotropic scaling, which are desirable for neural networks trained with gradient descent owing to their stochastic nature. These properties make CKA a precise and reliable metric for measuring similarities between neural networks, particularly when comparing potential models from malicious clients.
CKA has been utilized to address data heterogeneity issues, taking into account that the similarity of non-IID data is notably lower than that of IID data [20], particularly in specific layers [19]. However, CKA has yet to be leveraged as a defense mechanism against attacks involving models from malicious clients where the data are distributed across the clients in a non-IID manner.
III Threat Model
III-A Attacker’s Goal
We consider two primary attack strategies in the context of FL: untargeted model poisoning and targeted backdoor attacks. Untargeted attacks (Fang attacks [8]) aimed to evade robust aggregation rules like Krum and Coordinate-wise Median or Trimmed Mean. Attackers manipulate the model parameters to degrade the global model’s overall accuracy. Targeted backdoor attacks, represented by Bhagoji et al. [3], aim to deceive the global model into misclassifying specific data samples the attacker chose. The objective is to assign a different label chosen by the attacker while maintaining high accuracy for the remaining classes, making the attack more inconspicuous.
III-B Attackers’ Capabilities
  • An attacker on the client side can control multiple compromised clients.
  • An attacker has full control over at most $k<n/2$ clients.
  • An attacker has knowledge about compromised clients’ data samples, current local model parameters, global model parameters of the previous round, and hyperparameters, such as learning rate, optimization method, loss function, batch size, and the number of local epochs.
  • An attacker has no knowledge or control over the server and honest clients.
III-C Attack Strategy
III-C1 Untargeted Model Poisoning Attacks
These attacks based on [8] specifically aim to break Krum and Coordinate-wise Median (Coomed) aggregation rules, known to be byzantine failure tolerant. Their primary objective is to disrupt the model training process, thereby diminishing global test accuracy. The attacks involve manipulating model parameters, such as flipping the sign of malicious parameters, to steer the model in the opposite direction from its uninterrupted trajectory. Specifically, we denote the attack against Krum as Untargeted-Krum and the attack against Coomed as Untargeted-Med.
Untargeted-Krum [8] alters malicious parameters resembling benign ones to maximize the chances of being selected by Krum. Specifically, the optimization problem is as follows:
$\max_{\lambda}\lambda$
subject to $\text{subject to }w_{1}$ $=Krum(w_{1},...,w_{m},w_{m+1},...,w_{n})$
$w_{1}$ $=w_{G}-\lambda s$
$w_{i}$ $=w_{1},\text{ for }i=2,3,...,m$
where $m$ is the number of attackers, $n$ is the total number of selected clients, $w_{G}$ is the global model from the previous iteration, and $s$ is the sign of the global model parameter with no attack. The upper bound of the $\lambda$ is solved as follows:
$\lambda\leq$ $\frac{1}{(n-2m-1)\sqrt{d}}\cdot\min_{m+1\leq i\leq n}\sum_{l\in\Gamma_{w_{i}}^{n-m-2}}D(w_{l},w_{i})$
$+\frac{1}{\sqrt{d}}\cdot\max_{m+1\leq i\leq n}D(w_{i},w_{G})$
where $d$ is the number of parameters in the global model, $\Gamma$ is the set of $n-m-2$ benign local models having the smallest Euclidean distance to $w_{i}$, and $D$ is the Euclidean distance. $\lambda$ is halved until one of the compromised models is selected, or $\lambda$ is less than a threshold.
Untargeted-Med [8] manipulates the model parameters based on their maximum and minimum so that chosen coordinate-wise median values direct toward an inverse direction. The attack starts with defining $w_{max,j}$ and $w_{min,j}$, which indicates the maximum and minimum of the $j$th local model parameters on the benign clients, respectively.
$w_{max,j}$ $=\text{max}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$
$w_{min,j}$ $=\text{min}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$
Also, to avoid sampled $m$ numbers being outliers, if $s_{j}=-1$, $m$ numbers in [$w_{max,j},b\cdot w_{max,j}$] (when $w_{max,j}>0$) or [$w_{max,j},b/w_{max,j}$] (when $w_{max,j}\leq 0$) are randomly sampled. Otherwise, $m$ numbers [$w_{min,j}/b,w_{min,j}$] (when $w_{min,j}>0$) or [$b\cdot w_{min,j},w_{min,j}$] (when $w_{min,j}\leq 0$) are randomly sampled. Since the attack does not depend on $b$, we set $b=2$ as the same as [8].
III-C2 Targeted Backdoor Attack
This attack is based on [3]. Each malicious client owns one mislabeled image samples and trains the local model on it. We trained the model on backdoor tasks along with the main task(s); the training went on for both malicious and benign tasks to maintain benign accuracy such that side-job training (i.e., backdoor training) remains stealthy. Then, the attack boosted malicious clients’ updates to negate the combined effect of the benign client:
$w_{i}^{t}=w_{G}^{t-1}+\alpha_{m}(w_{i}^{t-1}-w_{G}^{t-1})\text{ for }i=1,...,m$
where $t$ is the current epoch and $\alpha_{m}$ is a boosting factor. In CIFAR10, for example, we manipulate the model parameters to misclassify an image of ’airplanes’ as ’bird’ whereas normally classifying the image of other classes like ’cat,’ ’horse,’ or ’truck.’
IV FedCC: Robust Aggregation Against Poisoning Attacks
Figure 2: An Overall Mechanism of our FedCC.
IV-A Overview of FedCC
Figure 2 depicts an overview of FedCC. (1) A server initializes and advertises a global model to $n$ participating clients, selected out of total $k$ clients with a probability of fraction $C$. (2) The clients individually train their local models using their respective datasets and send the trained weights to the server for aggregation. Our proposed method intervenes at this moment. (3) The server extracts PLRs from the global model and each local model, and (4) compares RBF Kernel CKA values between them. (5) The CKA values are then clustered into two, potentially representing benign and malicious clients. (6) Since the server cannot assure whether the clients with higher similarities are malicious, it performs a coordinate-wise average of the weights from clients belonging to the cluster with a larger number of members, backed by an assumption that the number of malicious clients does not exceed $n/2-1$. (7) Finally, the server updates the global model for the current epoch and proceeds to the next iteration by distributing the updated global model to newly selected $n$ clients.
IV-B Detailed Design
When filtering malicious clients in the presence of non-IID data, a major challenge is the server’s lack of knowledge regarding the data distribution. Specifically, even if the data is benign, local models can exhibit significant angular or magnitude differences if the data is non-IID. We thus require a similarity metric independent of data distribution and not influenced by the distance or direction of model parameters.
Figure 3: Cluster Distance of Each Layer
TABLE I: Comparison of Performance with Various Similarity Metrics
Method Fang-Med Fang-mKrum Targeted
IID NIID IID NIID IID NIID
Kernel CKA 69.20 41.00 70.22 43.24 71.44/6e-07 54.62/0.0118
Linear CKA 10.00 13.13 64.09 39.55 71.02/0.0007 49.53/0.0616
MMD 63.39 40.90 69.69 32.27 70.85/1e-09 50.51/9e-05
Cosine 68.82 33.90 68.81 10.04 69.76/0.0002 53.66/0.0529
Euclidean 69.06 27.82 68.54 41.57 69.17/0.0221 52.20/0.0015
Penultimate Layer is the output of the second-to-last layer before the softmax layer in CNN. In federated learning, where the server cannot access local training data, it receives trained model parameters (weights) from selected clients. Therefore, the collected weights are the only available information for determining whether a client is malicious. PLR takes the end result of the convolution and pooling process that will reach a classification decision. Thus, we assumed that computed weights of the second last layer might contain the essential information that strongly influences classification results.
To justify the utilization of PLRs, we measure the distance between two clusters of clients (benign and malicious) using dendrograms with a single linkage metric and correlation distance method. Figure 3 presents cluster distances of each layer under different attack scenarios, such as untargeted-mKrum attacks in IID or non-IID settings. Observably, the cluster distance is the most considerable in PLR to other layers, indicating the highest degree of differences. Based on this observation, we conclude that PLR contains the most distinguishable and indicative information. Therefore, we utilize the weights of the penultimate layer to compare their similarity to the global model.
Kernel CKA[13] is a similarity metric to assess the similarity between two neural networks. It compares representations at different layers or models trained in various ways. The sophisticated similarity comparison ability of Kernel CKA makes it a suitable choice for our purpose.
To justify the use of Kernel CKA, we compared its performance with other similarity metrics such as linear CKA, MMD (not normalized CKA), cosine (angle), and Euclidean similarity (distance). We evaluated the test accuracy of FedCC using the CIFAR10 dataset under two untargeted attacks and one targeted attack, separately in IID and non-IID settings. The results are summarized in Table I. In the ’targeted’ column, the values represent the combination of test accuracy and backdoor confidence, with higher test accuracy and lower backdoor confidence indicating better performance. We observed that Kernel CKA consistently yielded the highest performance across multiple experiments. Therefore, we decided to utilize Kernel CKA to compare PLR similarities.
Algorithm 1 FedCC
Input global_w, n local_w $w_{1}$, …, $w_{n}$
      Output agg_w
1:for $i<n$ do $\triangleright$ Get PLRs of local and global models
2:     local_plrs[$i$] $\leftarrow$ local_ws[second_last_layer]
3:end for
4:glob_plr $\leftarrow$ global_w[second_last_layer]
5:for $i<n$ do $\triangleright$ Apply kernel CKA to the plrs
6:     cka[$i$] $\leftarrow$ cka(glob_plr, local_plrs[$i$])
7:end $\triangleright$ Apply Kmeans clustering algorithm
8:kmeans = kmeans(n_cluster=2, cka[$i$])
9:labels = kmeans.labels_
10:count = counter(labels)
11:suspect $\leftarrow$ 1
12:if count[00] $\leq$ count[$1$then
13:     suspect $\leftarrow$ 0
14:end $\triangleright$ Define suspect based on the size of cluster
15:suspects = where(label == suspect)
16:for $i$ in range($n$do
17:     for $s$ in suspects do $\triangleright$ Set scale to 0 for the suspects
18:         scale[$s$] $\leftarrow 0$
19:     end for
20:     if scale[$i$] != 0 then
21:         scale[$i$] $\leftarrow 1$
22:         selected_param.append(scale[$i$] * local_ws)
23:     end if
24:end $\triangleright$ Average the selected parameters
25:agg_weights = mean(scale*selected_param)
26:return agg_weights
FedCC is an aggregation method that combines selected weights without scaling, based on the measurement of Kernel CKA similarities between the global and local models’ PLRs. A pseudo algorithm is described in Algorithm 1. A server first extracts PLRs of the global and local models and calculates the kernel CKA between them. We specifically focus on comparing PLRs because they exhibit the most distinguishable differences among differently trained models, as demonstrated in previous sections. Additionally, we utilize RBF Kernel CKA, as it effectively captures similarity differences between non-IID and malicious clients. The similarity values obtained are then clustered using a simple K-means clustering algorithm. Any clustering method that produces only two clusters can be used. Considering the assumption that the number of attackers is less than half of the total number of clients ($n/2$), we designate the clients in the cluster with fewer members as suspects. The indices of these suspicious clients are stored in a list called ‘suspects.’ We proceed by selectively averaging the weights of clients whose index does not belong to the ‘suspects’ list, excluding the weights of the suspicious clients from the aggregation. Note that the weights are averaged without scaling, as scaling could impede the model from convergence even when the loss reaches optima. We named our proposed method FedCC.
Compared to other methods such as Krum, Bulyan, and FLARE, which either require prior knowledge of the number of attackers or allow the server to have access to raw data, FedCC does not rely on any information about the clients or any instances of data. This ensures strong privacy preservation within the federated learning framework. Furthermore, FedCC effectively distinguishes between malicious and benign clients, even in practical but challenging scenarios where benign clients have non-IID data.
V Experiments
V-A Dataset
We use three benchmark vision datasets: Fashion-MNIST (fMNIST) [24], CIFAR10, and CIFAR100 [14].
fMNIST consists of 60,000 training and 10,000 testing samples of 28$\times$28 grayscale images and associating labels of 10 classes. The labels include T-shirt/top, Trouser, Pullover, Dress, Coat, Sandal, Shirt, Sneaker, Bag, and Ankle boot.
CIFAR10 consists of 50,000 training and 10,000 testing examples of 32$\times$32 color images and associating labels of 10 classes. The classes include airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck.
CIFAR100 is the same as CIFAR10 except for the categorization. It is labeled as 100 classes, each containing 600 images. Thus, there are 500 training and 100 testing images per class. Each class (i.e., ‘fine’ label) belongs to one of the 20 superclasses.
V-B Non-IID Simulation
Non-IID is a more feasible assumption considering the diverse and massive nature of data in practice. To standardize non-IID settings, we use Dirichlet distribution, which is well-known for its capability of representing real-world data [15]. A concentration parameter $\alpha$ controls the degree of non-IID; the smaller $\alpha$ is, the more likely the clients hold examples from only one randomly chosen class.
V-C Model Architecture
Considering that the end devices are normally incapable of handling heavy computation due to resource or communication constraints, we used lightweight CNN architectures for the experiments. For fMNIST, we used a 4-layer-CNN with two convolutions and two fully connected (FC) layers, incorporating dropout with probabilities of 0.25 and 0.5 before and after the FC layer, respectively. For CIFAR10, we utilized a 5-layer-CNN with three convolutions, each followed by a max-pool layer and two FC layers, applying dropout with a probability of 0.5 before the first FC layer. Finally, for CIFAR100, we adopted the LeNet architecture. Table VI, Table VII, and Table VIII in Appendix summarize model architectures for each benchmark dataset.
V-D Setup
We implement two untargeted attacks in [8] and one targeted attack [3], defense baselines, including Krum [4], Coomed [26], Bulyan [11], and FLARE [22], and FedCC. We have ten clients with a client participation fraction of 1.0 unless stated otherwise. The number of attackers is 30% of the total clients (three) in untargeted and 10% (one) in targeted attacks as in [22]. We also assumed all clients have the same NN architecture. Under the targeted backdoor attack, benign clients go through three, whereas compromised clients undergo six local epochs. We train each client’s local model using an Adam optimizer with a learning rate of 0.001, following the experiment in [22]. All results are averaged after running each experiment three times.
TABLE II: Accuracy of no-attack scenario, boldface and underlining refer to the highest and the second highest accuracy, respectively.
Scenarios Dataset FedAvg Krum Coomed Bulyan FLARE FedCC
No Attack non-IID fMNIST 69.68 43.10 76.12 68.78 42.82 71.92
CIFAR10 53.06 16.16 45.55 46.67 27.33 50.80
CIFAR100 17.22 7.27 12.82 15.21 0.99 17.62
No Attack IID fMNIST 89.89 83.81 88.91 89.49 76.86 89.71
CIFAR10 69.17 56.56 67.34 67.81 46.84 69.03
CIFAR100 21.49 7.73 15.56 20.55 7.59 22.05
V-E Evaluation Metrics
We assessed the performance of our defense mechanism using two metrics: backdoor confidence (confidence) and global model test accuracy (accuracy). The confidence indicates the likelihood of misclassification, with values ranging from 0 to 1 (lower values indicating better defense performance). Accuracy measures the overall test accuracy of the global model when various defense methods are applied when attacks exist, with values ranging from 0 to 100 (higher values indicating better performance). We reported both metrics in the presence of backdoor attacks while focusing solely on accuracy for untargeted attacks—as the attacker’s goal is to deteriorate the overall accuracy.
VI Results and Discussion
In Table II, which represents the baseline experiments without any attacks, we can observe the accuracy of different methods in both non-IID and IID settings. FedCC achieves accuracy comparable to that of the best-performing method, FedAvg. The slightly higher accuracy of FedAvg than FedCC may be attributed to its wholesome averaging, which leads to a more generalized model when there is no attack. Nonetheless, FedCC surpasses all other baseline methods with CIFAR100, which has more classes, making it harder to achieve better accuracy in FL[7].
TABLE III: Accuracy under untargeted and targeted attacks in Non-IID (top) and IID setting (bottom)
Scenarios Dataset FedAvg Krum Coomed Bulyan FLARE FedCC
Fang-Krum non-IID fMNIST 64.63 16.51 57.12 13.39 49.54 66.27
CIFAR10 33.69 15.38 35.70 19.23 17.03 37.40
CIFAR100 2.27 1.00 4.85 0.98 7.46 14.51
Fang-Med non-IID fMNIST 16.32 49.33 68.90 64.12 52.25 72.76
CIFAR10 10.02 25.06 40.23 32.47 14.59 42.37
CIFAR100 1.00 6.24 10.27 6.91 1.00 16.12
Targeted non-IID fMNIST 79.07 43.46 71.25 73.58 60.57 84.54
CIFAR10 30.28 16.05 28.87 25.77 10.52 36.20
CIFAR100 9.84 7.24 11.02 12.01 1.04 13.81
Fang-Krum IID fMNIST 75.55 31.66 87.62 50.30 79.16 89.23
CIFAR10 49.67 40.86 57.40 12.67 25.77 67.12
CIFAR100 13.72 1.04 6.17 1.59 7.49 18.47
Fang-Med IID fMNIST 20.86 85.33 86.70 87.45 71.08 88.44
CIFAR10 9.51 54.28 59.20 57.69 49.82 64.74
CIFAR100 0.87 12.27 14.43 12.56 5.83 17.83
Targeted IID fMNIST 89.32 85.52 88.45 89.35 67.65 90.16
CIFAR10 52.20 45.37 49.27 55.77 10.09 51.57
CIFAR100 12.73 8.37 15.91 16.59 1.71 18.08
VI-A Non-IID Data Environment
Following the aforementioned non-IID simulation, we set $\alpha$ to 0.2 to create a disjoint client training data and distribute each set of data to each client.
VI-A1 Untargeted Model Poisoning Attacks
The upper section of Table III provides the global model’s accuracy across three datasets under two untargeted (Fang-Krum and Fang-Med) and Targeted attacks in a non-IID environment. One notable finding is that robust aggregation algorithms often yield lower accuracy compared to simple averaging (no defense). This discrepancy arises from the imperfect identification of malicious clients, leading to the erroneous aggregation of their weights along with benign clients. When applying Krum and Bulyan against Fang-Krum, the accuracy drops by a substantial 59.30% and 65.68%, respectively, compared to FedAvg. This decline can be attributed to Fang-Krum’s intentional design to degrade accuracy against Krum. Additionally, Coomed struggles to effectively mitigate this attack as the median values of weights often fail to represent benign clients, particularly in non-IID settings, accurately. Furthermore, the relatively low accuracy of FLARE is due to weight scaling, even when the model already exhibits satisfactory performance. In contrast, FedCC, which selectively averages weights, stands out by neither boosting nor minimizing weights, resulting in higher performance than other methods. Notably, FedCC achieves the highest accuracy (66.27%, 37.40%, and 14.51% for fMNIST, CIFAR10, and CIFAR100, respectively), with a significant leap in experiments with CIFAR100.
Turning our attention to the Fang-Med attack, the accuracy of FedAvg experiences a substantial drop due to the attack’s tendency to deviate from median values, creating outliers. This behavior explains the relatively high accuracy when Coomed and Bulyan are applied because coordinate-wise median and trimmed mean methods disregard outliers. Meanwhile, FLARE performs poorly for a reason similar to the one mentioned above. For both attacks, we can observe that Krum, which selects a single local model as the global model, is less robust than Coomed, which aggregates multiple local models. Similarly, Bulyan, which takes averages of multiple chosen clients, exhibits less robustness than Coomed, particularly in non-IID data environments. Notably, FedCC achieves the highest accuracy (72.76%, 42.37%, and 16.12% for each dataset) without sharing raw data with the server or other clients with non-IID data. This exceptional performance of FedCC can be attributed to two factors: 1) highly distinguishable information within PLRs, and 2) higher CKA similarity between benign clients than the one between benign and malicious clients.
VI-A2 Targeted Backdoor Attacks
Figure 4: Confidence of Backdoor Task in a Non-IID Setting (top) and an IID setting (bottom)
The fourth row of Table III summarizes accuracy under targeted backdoor attacks in a non-IID setting. To further illustrate the impact of these attacks, we include a visual representation of the backdoor confidence in the first row of Figure 4.
It becomes evident that Krum, which selects a single client’s weights as the global model, exhibits diminished robustness compared to other methods that aggregate the weights of multiple clients; coordinate-wise median (Coomed) and trimmed mean with multi-Krum (Bulyan) demonstrate relatively higher test accuracy. Conversely, FLARE decreases the main task test accuracy for similar reasons observed in the experiments mentioned above. Nonetheless, FedCC achieves the highest test accuracy. This outcome underscores the superiority of filtering malicious clients through CKA similarity comparison, surpassing the effectiveness of first-order statistical methods like mean or median.
Notably, Figure 4 illustrates how FedCC significantly nullifies backdoor confidence. Compared to oscillating values of other methods, the stable low confidence values further emphasize FedCC’s resilience and independence from the specific data distribution across clients. This advantage stems from its effective utilization of CKA, enabling the calculation of similarity values even when models are trained on different datasets with varying data distributions. Whereas it is true that other baselines, such as Bulyan, also reduce confidence, we argue that FedCC not only preserves but potentially increases accuracy by precisely discriminating against malicious clients.
VI-B IID Data Environment
In an IID setting, we evenly divide the training dataset among all clients so that each client’s data distribution is identical and of the same size.
VI-B1 Untargeted Model Poisoning Attacks
The bottom section of Table III (fifth to seventh rows) provides valuable insights into the test accuracy of a global model trained on three datasets under untargeted attacks in IID settings. Although the trends observed in non-IID experiments appear similar, the accuracy values are slightly higher in this case due to the identical nature of the data.
Under the Fang-Krum attack, we can observe relatively low accuracy values for Krum and Bulyan, as the attack explicitly targets Krum and its variations. However, when Coomed is applied, the accuracy increases from 75.55% (FedAvg) to 79.16% (Coomed), indicating that coordinate-wise median values closely align with the weights of the global model with minimal deviation. FLARE’s low accuracy is due to the continuous scaling of weights even when the model is already optimized. In contrast, FedCC does not scale weights but focuses solely on optimizing the model to the minimum by averaging chosen clients. As a result, FedCC achieves the highest accuracy (89.23%, 67.12%, and 14.51% for each dataset) by accurately identifying and filtering out compromised clients.
Under the Fang-Med attack, the accuracy of FedAvg experiences a significant drop (89.89% to 20.86%) due to the peculiarity of outliers when the clients have IID data. Notably, Krum, Coomed, and Bulyan demonstrate similar accuracy values, indicating that coordinate-wise median values indeed represent benign clients’ parameters. Meanwhile, FedCC consistently achieves the highest accuracy, with values of 88.44%, 64.74%, and 17.83% for each dataset. Notably, the remarkable growth observed in experiments with CIFAR100, with a 66% and 58.7% increment under Fang-Krum and Fang-Med attacks, respectively, surpasses the second-highest improvement. These results strongly validate the effectiveness of FedCC in mitigating untargeted model poisoning attacks in a simple IID setting.
VI-B2 Targeted Backdoor Attacks
The last row of Table III and the bottom three figures of Figure 4 are the test accuracy and backdoor confidence under targeted backdoor attacks in an IID setting. In targeted backdoor attacks, the global model is covertly trained to misclassify a specific attacker-chosen class to an attacker-chosen label while leaving the remaining classes intact. As a result, the test accuracy has remained relatively high across the different methods.
However, a critical observation is the remarkable confidence reduction achieved by FedCC. The backdoor confidence with FedCC approaches zero, effectively neutralizing the backdoor task. In comparison, the attack confidence of FedAvg and Bulyan, despite exhibiting higher test accuracy for CIFAR10, remains higher than that of FedCC. Note that the confidence of the other methods fluctuates by a large margin. Specifically, FLARE’s backdoor confidence oscillates even though it occasionally records lower confidence than ours (e.g., CIFAR10). This means that its performance is less stable than that of FedCC, which consistently minimizes backdoor confidence. It is crucial to emphasize that reducing attack confidence is as significant as maintaining the main test accuracy since the attacker’s objective is to train the model on backdoor tasks stealthily, but the defender’s goal is to nullify them. Hence, our proposed method achieves the highest accuracy for fMNIST and CIFAR100 and comparable accuracy for CIFAR10 while effectively nullifying the backdoor confidence. This outcome underscores the effectiveness and robustness of our approach in mitigating targeted backdoor attacks.
VII Performance in Varying FL Settings
We additionally measure the effectiveness of FedCC in various non-IID settings, including different numbers of malicious clients, fractions of participation, and numbers of local epochs. Since Coomed was as effective as ours, and FLARE also used PLRs, we compare the three methods to FedCC, including FedAvg.
VII-A Impact of the Number of Malicious Clients
Figure 5: Accuracy with different numbers of malicious clients
We assess the impact of different numbers of malicious clients under untargeted attacks, specifically Fang-Krum and Fang-Med. In this set of experiments, we fix the total number of clients and participation fraction at 10 and 1, respectively. To adhere to the assumption that the number of attackers is less than half of the participating clients, we investigate the test accuracy considering scenarios involving up to four malicious clients.
Figure 5 illustrates the accuracy in the given experimental settings, and it is evident that FedCC outperforms the other defense methods. We observe a general trend where the accuracy of defense methods decreases as the number of attackers increases. Notably, Coomed experiences the most significant drop in accuracy when transitioning from one to four malicious clients. This can be attributed to the fact that as the participation of malicious clients increases, there is a higher likelihood that the median values will be influenced by the malicious contributions.
In contrast, FedCC does not rely solely on geometric measures such as angles or distances but instead leverages hidden correlations between networks to identify and filter out malicious clients. As a result, FedCC demonstrates superior accuracy in filtering out malicious clients, regardless of the number of attackers involved. It is important to note that the accuracy degradation observed as the number of malicious clients increases primarily due to aggregating fewer clients rather than the malicious clients themselves being selected. These findings underscore the robustness and effectiveness of FedCC in defending against untargeted attacks, as it consistently outperforms other defenses across varying numbers of malicious clients.
VII-B Impact of Fraction
TABLE IV: Accuracy under untargeted attacks with different fractions of clients being selected. fM refers to fMNIST, C10 refers to CIFAR10, Med refers to Coordinate wise median.
Fang-Krum Fang-Med
Frac Data FedAvg Med FLARE FedCC FedAvg Med FLARE FedCC
0.1 fM 55.31 49.83 34.02 64.83 16.57 66.41 52.24 69.52
C10 10.06 22.55 14.50 20.49 10.00 15.33 10.00 29.81
0.3 fM 64.22 57.52 10.00 73.55 16.26 58.07 10.00 61.12
C10 24.24 12.59 10.00 27.81 10.98 22.61 10.00 38.27
0.5 fM 62.37 58.20 10.00 76.41 18.36 62.49 10.00 69.05
C10 23.99 17.28 10.06 34.32 9.87 27.83 10.00 37.27
We examine the impact of the participating fraction of clients where the number of total clients is fixed at 100. The fractions are 0.1, 0.3, and 0.5, such that the numbers of clients being aggregated are 10, 30, and 50, respectively. The proportion of malicious clients is fixed at 0.3, such that the malicious clients are 3, 9, and 15, respectively. Table IV summarizes the accuracy. We can observe that the accuracy tends to increase as the participation fraction is greater under Fang-Krum attacks. It implies that the more client participation in a non-IID setting, the more accurate the global model is if the attacks are mitigated. Under Fang-Med attacks, the accuracy is the highest with FedCC and second highest with Coomed. Throughout the experiment, FLARE prevented the global model from convergence, and the test accuracy fluctuated. The fluctuation is presumably due to the weight scaling while randomly selecting clients. Distinct from previous experiments, the server chooses a fraction of the client every round, meaning not-yet-trained models could be selected. In this environment, weight scaling leveraged by FLARE constantly improperly scales the local weights, causing divergence and fluctuation. These experimental results indicate that selectively choosing clients based on CKA similarity is more reliable than taking the median value of each coordinate in a non-IID setting. We also observed that as the participation fraction grows, the accuracy grows. This is due to the rising number of aggregated local models; the more models are aggregated, the more general global model is generated.
VII-C Impact of the Number of Local Epochs
TABLE V: Accuracy under Backdoor Attack with doubled numbers of local epochs
Targeted (Diff Local Epoch)
Dataset FedAvg Coomed FLARE FedCC
fMNIST 69.40 72.11 72.32 82.31
CIFAR10 53.85 49.18 21.61 54.94
Figure 6: Backdoor confidence when the numbers of local epochs are doubled.
To examine the impact of the number of local epochs, we doubled up the number of local epochs such that the local models are trained for six epochs in benign clients and twelve epochs in malicious clients. The main task accuracy and confidence are summarized and illustrated in Table V and Figure 6. As expected, increasing the number of local epochs would not affect the accuracy or confidence but bring the convergence forward while being better at maintaining test accuracy. It is true that Coomed and FLARE significantly reduce the backdoor confidence, as well as FedCC; however, the test accuracy has even increased with FedCC due to its precise discrimination against malicious clients.
VIII Conclusion and Future Work
FL has emerged in response to the rising concerns about privacy breaches while employing AI techniques. FL, however, is vulnerable to model poisoning attacks due to a server’s blindness to local datasets, making it challenging to assume data distribution and model parameter integrity. In line with this, existing robust aggregation algorithms and defense approaches are orthogonal to various attacks and lack consideration of non-IIDness. Throughout exhaustive experiments, FedCC mitigates both untargeted and targeted (or backdoor) attacks while demonstrating its effectiveness in non-IID data environments. We leave theoretical guarantees of FedCC and experiments when clients have distinct neural network architectures or tasks as future work.
Appendix
VIII-A Dendrograms to Measure Cluster Distance
Figure 7: Dendrogram in a non-IID setting
Figure 7 illustrates the distance between two clusters of clients using dendrograms with a single linkage metric and correlation distance method. The first two and last columns correspond to untargeted (Fang-Krum, Fang-Med) and targeted attacks, respectively. The total number of clients is set to 10; the participation fraction is 1.0. The number of malicious clients is 3 for untargeted and 1 for a targeted attack, with fedAvg. Observably, the height difference between the root of clusters is maximum in the second last layer (PLR).
VIII-B Model Architectures
The tables below summarize model architectures for each benchmark dataset (fMNIST, CIFAR10, CIFAR100). Considering that the end devices are generally incapable of handling heavy computation due to resource or communication constraints, we used lightweight models for the experiments.
TABLE VI: CNN Architecture for fMNIST dataset
Layer In Out Ker / Str / Pad Activation
conv2d_1 1 64 $5\times 5$ / 1 / 0 ReLU
conv2d_2 64 64 $5\times 5$ / 1 / 0 ReLU
dropout - - 0.25 -
flatten - 25600 - -
fc_1 25600 128 - -
dropout - - 0.5 -
fc_2 128 10 - -
TABLE VII: CNN Architecture for CIFAR10 dataset
Layer In Out Ker / Str / Pad Activation
conv2d_1 3 64 $3\times 3$ / 1 / 0 ReLU
maxpool2d $2\times 2$ / - / 0 -
conv2d_2 64 64 $3\times 3$ / 1 / 0 ReLU
maxpool2d $2\times 2$ / - / 0 -
conv2d_3 64 64 $3\times 3$ / 1 / 0 ReLU
maxpool2d $2\times 2$ / - / 0 -
flatten - 256 - -
dropout - - 0.5 -
fc_1 256 128 - -
fc_2 128 10 - -
TABLE VIII: LeNet Architecture for CIFAR100 dataset
Layer In Out Ker / Str / Pad Activation
conv2d_1 3 6 $5\times 5$ / 1 / 0 ReLU
maxpool2d $2\times 2$ / 2 / 0 -
conv2d_2 6 16 $5\times 5$ / 1 / 2 ReLU
maxpool2d $2\times 2$ / 2 / 0 -
flatten - $4$
fc_1 44944 120 - ReLU
fc_2 120 84 - ReLU
fc_3 84 100 - -
Acknowledgment

### v3

@@
-finantial support acknowledgement
+FedCC: Robust Federated Learning against Model Poisoning Attacks
 Abstract
-Federated Learning, designed to address privacy concerns in learning models, introduces a new distributed paradigm that safeguards data privacy but differentiates the attack surface due to the server’s inaccessibility to local datasets and the change in protection objective–parameters’ integrity.
-Existing approaches, including robust aggregation algorithms, fail to effectively filter out malicious clients, especially those with non-Independently and Identically Distributed data.
-Furthermore, these approaches often tackle non-IID data and poisoning attacks separately.
-To address both challenges simultaneously, we present FedCC, a simple yet novel algorithm.
-It leverages the Centered Kernel Alignment similarity of Penultimate Layer Representations for clustering, allowing it to identify and filter out malicious clients by selectively averaging chosen parameters, even in non-IID data settings.
-Our extensive experiments demonstrate the effectiveness of FedCC in mitigating untargeted model poisoning and backdoor attacks.
-FedCC reduces the attack confidence to a consistent zero compared to existing outlier detection-based and first-order statistics-based methods.
+Federated learning is a distributed framework designed to address privacy concerns.
+However, it introduces new attack surfaces, which are especially prone when data is non-Independently and Identically Distributed.
+Existing approaches fail to effectively mitigate the malicious influence in this setting; previous approaches often tackle non-IID data and poisoning attacks separately.
+To address both challenges simultaneously, we present FedCC, a simple yet effective novel defense algorithm against model poisoning attacks.
+It leverages the Centered Kernel Alignment similarity of Penultimate Layer Representations for clustering, allowing the identification and filtration of malicious clients, even in non-IID data settings.
+The penultimate layer representations are meaningful since the later layers are more sensitive to local data distributions, which allows better detection of malicious clients.
+The sophisticated utilization of layer-wise Centered Kernel Alignment similarity allows attack mitigation while leveraging useful knowledge obtained.
+Our extensive experiments demonstrate the effectiveness of FedCC in mitigating both untargeted model poisoning and targeted backdoor attacks.
+Compared to existing outlier detection-based and first-order statistics-based methods, FedCC consistently reduces attack confidence to zero.
 Specifically, it significantly minimizes the average degradation of global performance by 65.5%.
-We believe that this new perspective of assessing learning models makes it a valuable contribution to the field of FL model security and privacy.
-The code will be made available upon paper acceptance.
-Index Terms:
-Backdoor Attack, Defense, Federated Learning, Model Poisoning Attack, Non-IID, Robust Aggregation
-[…]
-=-21pt
-I Introduction
-Federated Learning (FL) [17] has emerged as a promising privacy-preserving model training approach in response to growing concerns about privacy breaches and data leakage caused by centralized learning.
-While Machine Learning (ML) and Deep Learning (DL) algorithms are popular for their capability, such as personalized recommendations and accurate forecasts, their centralized nature risks data privacy.
-FL allows data to remain on client devices but exchanges their local model parameters only, avoiding the need for central data collection.
-FedAvg [17] is the first algorithm run on the server to aggregate clients’ model parameters.
-It involves a server initializing and advertising a global model to all clients, who then train the model with their private data and upload the computed parameters to the server.
-The server averages these parameters to update the global model.
-These steps are repeated until a stopping criterion is met.
-FL is especially suitable for data-sensitive environments since raw data remains on clients, reducing the risk of data leakage on the server or in-between communication.
-However, due to its distributed nature, FL is susceptible to model poisoning attacks [12].
-The server cannot directly examine local datasets’ data quality or model parameter integrity, allowing compromised clients or attackers to manipulate local models.
-The attacks can degrade global model performance indiscriminately (untargeted attacks) [8, 2, 18] or cause incorrect predictions on specific inputs (targeted attacks) [21, 23, 3, 25].
-Backdoor attacks, a type of targeted attack, are stealthier by maintaining overall performance while misclassifying specific inputs [1].
-Some preliminary defenses propose robust aggregation methods, but they may compromise privacy by sharing raw data or exposing data distribution to unreliable parties [6, 22].
-Existing defenses often target IID data settings, which are rare in practice.
-Non-IID data, with variations in class distribution, features, or labels across clients [12], can gradually degrade model performance and complicate the detection of malicious clients [28, 27], as the impact of attacks increases with the degree of non-IID [16, 18].
-In this study, we present FedCC, a novel defense mechanism for FL that mitigates untargeted model poisoning and targeted backdoor attacks, even with non-IID data.
-FedCC leverages the Centered Kernel Alignment (CKA) of Penultimate Layer Representations (PLRs) to distinguish between malicious and benign clients while maintaining data privacy by avoiding sharing raw data or additional information between the server and clients.
-PLRs contain highly distinguishable features, maximizing the differences in CKA scores between benign and malicious clients, even with non-IID data and no server knowledge of local data.
-Empirical experiments show that FedCC outperforms existing defenses that rely on first-order statistics.
-It significantly improves global accuracy against untargeted attacks (65.5% compared to similar approaches) and reduces attack confidence in targeted backdoor attacks to almost zero while preserving main task accuracy.
+We believe that this new perspective on aggregation makes it a valuable contribution to the field of FL model security and privacy.
+The code will be made available upon acceptance.
+Keywords:
+Federated learning model poisoning attack backdoor attack robust aggregation
+1 Introduction
+Federated Learning (FL) [22] is a distributed model training framework designed to preserve privacy.
+FL allows data to remain on client devices but exchanges their local model parameters only for model training.
+It thus reduces (data) privacy risks associated with centralized learning, not requiring central data collection.
+Thus, FL is especially suitable for data-sensitive environments since raw data remains on clients, reducing the risk of data leakage on the server or in between communication.
+However, its distributed nature makes FL susceptible to model poisoning attacks [12].
+The server cannot directly examine local datasets’ data quality or model parameter integrity, leaving compromised clients or attackers to manipulate local models.
+This can degrade global model performance indiscriminately (untargeted attacks) [7, 2, 26] or cause incorrect predictions on specific inputs (targeted attacks) [29, 31, 3, 33].
+Backdoor attacks are stealthier targeted attacks that maintain overall performance while misclassifying specific inputs [1].
+Some preliminary defenses propose robust aggregation methods, but they may compromise privacy by sharing raw data or exposing data distribution to untrusted parties [6, 30].
+While some defenses exist, most focus on IID data settings, which are rare in real-world scenarios.
+Non-IID data, with variations in class distribution and features across clients [12], can gradually degrade model performance and complicate the detection of malicious clients [39, 35], as the impact of attacks increases with the degree of non-IID [19, 26].
+In this study, we introduce FedCC, a defense mechanism that mitigates untargeted model poisoning and targeted backdoor attacks in both IID and non-IID settings.
+FedCC leverages the Centered Kernel Alignment (CKA) of Penultimate Layer Representations (PLRs) to distinguish malicious clients from benign ones.
+PLRs are known to be highly specific to local data distributions [36], making clients distinguishable, especially when data is non-IID.
+Exploiting CKA’s well-known capability of measuring similarity in high-dimensional space, such as neural network representations, we accurately measure the similarity between the PLR of different clients, further allowing more accurate detection of malicious clients.
+Empirical evidence supporting our rationale is shown in Figure 2 and Table 1.
+The use of PLRs shows the highest separability between benign and malicious clients in all settings, as seen by the largest distance between clusters in the PLR.
+The use of CKA, specifically kernel CKA, is justified by its highest performance in both targeted and untargeted IID and non-IID settings.
+Furthermore, our method maintains data privacy by using similarity measures between local and global models, not having shared data.
+We named our proposed method FedCC, Federated learning using Clustered CKA.
+Empirical experiments show that FedCC outperforms existing defenses that rely on first-order statistics, which usually suffer from a high false negative detection rate.
+It significantly improves global accuracy against untargeted attacks (66.27% compared to similar approaches) and reduces attack confidence in targeted backdoor attacks to almost zero while preserving main task accuracy.
 These results highlight FedCC’s effectiveness and robustness in mitigating attacks in FL settings, particularly in non-IID data settings.
 Our contributions are summarized as follows:
-  • We propose FedCC, a simple yet novel detection and aggregation method for FL, to defend against model poisoning attacks (untargeted model poisoning and targeted backdoor attacks) by comparing the CKA of PLRs of local and global models and averaging the selected local model parameters.
-  • We justify that PLR is the most distinguishable and indicative feature to classify malicious and benign clients by measuring the distance between two clusters of clients.
-  • We empirically show that CKA is the most sensitive and accurate metric to measure similarities between differently trained models, especially when data are non-IID, and the server does not have any knowledge about clients or any centralized data.
-  • We empirically demonstrated the effectiveness of FedCC for defending against model poisoning attacks.
-Our experimental results show that comparing the CKA of PLRs is more reliable in filtering malicious clients out than using first-order statistics.
-FedCC outperforms existing defenses even when data are non-IID.
-II Backgrounds and Related Works
-II-A Poisoning Attacks in FL
-FL is vulnerable to poisoning attacks due to its distributed nature and the server’s inability to examine dataset quality and model parameter integrity directly.
-Model poisoning is particularly destructive among various poisoning attacks, as adversaries stealthily manipulate local model parameters to maximize overall performance damage [3, 5].
-Thus, this paper focuses on model poisoning attacks.
-Model poisoning attacks allow client-side adversaries to alter local model parameters before sending them to the server.
-To avoid the global model from divergence, attackers optimize the compromised local model for both training loss and an adversarial objective, enhancing its stealthiness [5, 8].
-They can also adjust model hyperparameters, such as the learning rate, local epochs, batch size, and optimization objective, to manipulate training rules before and during local training.
-Based on the attacker’s goals, model poisoning attacks can be classified as untargeted or targeted.
-Untargeted attacks aim to degrade overall global test accuracy by manipulating compromised local models.
-In contrast, targeted attacks aim to mislead the model into misclassifying specific inputs, leaving other classes unaffected.
-This paper introduces a backdoor ability into targeted model poisoning attacks, where inputs belonging to a chosen class are intentionally misclassified to a chosen label, whereas the performance of other classes remains unaffected.
-Its added stealthiness makes the attack more challenging to detect [23, 21].
-II-B Robust Aggregation in FL
-Several Byzantine-robust aggregations based on summary statistics or anomaly detection have been proposed to mitigate model poisoning and backdoor attacks in FL.
-For example, Krum [4] selected the update with the smallest Euclidean distance to the remaining updates as the new global model.
-Still, it ignored outliers and assumed IID data only [2].
-Median [26] selects the coordinate-wise median of updates, also ignoring outliers.
-Trimmed Mean [26] excludes extreme values before averaging, and Bulyan [11] combines Krum and Trimmed Mean for selection.
-These methods rely on knowing the number of attackers, which is often unavailable in practice.
-Inspired by the fact that Sybil’s crafted model parameter’s directions are more similar to each other than to benign ones, Foolsgold [9] identified Sybils by measuring cosine similarity between local model gradients, marking the most similar ones as Sybils.
-However, it struggled with a single malicious client when there were no other malicious clients for comparison.
-Additionally, this approach may overfit non-IID data, ironically leading to suboptimal performance with IID data.
-FLTrust [6] used trust scores based on ReLU-clipped cosine similarities to weigh local clients’ parameters differently but risked privacy by requiring clients to share raw data with the server.
-Furthermore, precise weight scaling is paramount to ensure model convergence.
-Previous works often used summary statistics (mean or median), direction (cosine similarity), or distance (Euclidean) of weight vectors, but they struggled to effectively distinguish between benign clients with non-IID data and malicious clients, often misclassifying benign clients as malicious.
-The lack of performance might be partially attributed to the inadequacy of Euclidean geometry as a metric to compare neural network representations.
-In contrast, our proposed approach utilizes kernel-based metric (Kernel CKA) to measure the similarity between global and local models, reliably identifying compromised model parameters among the mixture of clients and enhancing defense against model poisoning attacks in FL.
-II-C Penultimate Layer Representation in FL
-Figure 1: CNN architecture and Penultimate Layer
-The penultimate layer is the second-to-last layer (before the softmax layer) of a neural network model, as illustrated in Figure 1.
+  • We propose FedCC, a novel and scalable defense method against model poisoning attacks using the CKA of PLRs of local and global models and performing layer-wise weighted aggregation of their model parameters.
+  • We justify that PLR is a highly distinguishable feature to detect malicious clients by measuring the discrepancy between two clusters of clients.
+  • We demonstrate that CKA provides the most sensitive and accurate similarity measure between differently trained models when the server has no knowledge about clients or centralized data, even in non-IID settings.
+  • We empirically demonstrate the effectiveness of FedCC through exhaustive experiments.
+Our results show that comparing the CKA of PLRs filters malicious clients more reliably than the first-order statistics.
+FedCC outperforms existing defenses even in non-IID settings.
+2 Backgrounds and Related Works
+2.1 Poisoning Attacks in FL
+FL is vulnerable to poisoning attacks due to its distributed nature, where the server cannot directly inspect dataset quality or model integrity.
+Model poisoning attacks are particularly destructive, as adversaries can stealthily manipulate local model parameters to degrade global model performance [3, 5], and this paper focuses on such attacks.
+In model poisoning attacks, client-side adversaries alter local model parameters before submitting them to the server.
+To avoid detection and prevent global model divergence, attackers optimize the local models for both training loss and an adversarial objective [5, 7].
+For instance, A3FL [37] dynamically fine-tunes backdoor triggers to make them harder to detect, while IBA [24] generates robust triggers using a generative network, exploiting the global model as a discriminator.
+Attackers can further manipulate hyperparameters, such as learning rate, local epochs, batch size, and regularization [17, 21], dynamically before and during local training to evade detection.
+In a similar vein, DBA [33] embedded split triggers into local training data, associating them with targeted incorrect predictions.
+Unlike centralized backdoor attacks, DBAs distribute malicious updates across multiple adversarial participants, making them harder to detect with anomaly detection techniques.
+Model poisoning attacks can be classified into untargeted and targeted types.
+Untargeted attacks degrade global model accuracy, while targeted attacks mislead the model to misclassify specific inputs without affecting other classes.
+Our defense mechanism mitigates both untargeted attacks by restoring global model accuracy and targeted backdoor attacks by overcoming the stealthiness and detection difficulty [31, 29].
+2.2 Robust Aggregation algorithm in FL
+Several Byzantine-robust aggregation methods, based on summary statistics or anomaly detection, have been proposed to mitigate model poisoning and backdoor attacks in FL.
+Krum [4], for instance, selects the update with the smallest Euclidean distance to others but assumes IID data and overlooks outliers [2].
+Multi-Krum, which averages updates from multiple clients, is known to be more effective in non-IID settings [4].
+Similarly, Median [34] computes the coordinate-wise median, Trimmed Mean [34] excludes extreme values, and Bulyan [10] combines Krum and Trimmed Mean for added robustness.
+These methods, however, require knowledge of the number of attackers, often unavailable in practice.
+Inspired by the observation that malicious clients’ model parameters exhibit higher similarity, Foolsgold [8] identified Sybils by measuring cosine similarity between gradients.
+While effective against multiple Sybils, it struggled when only a single malicious client exists or with IID data, where it can overfit.
+FLTrust [6] addressed this by assigning trust scores based on ReLU-clipped cosine similarities but risking privacy by raw data sharing.
+Similarly, Lockdown [11] improved robustness by pruning parameters unused by the majority of clients during training, while FLIP [38] rejected low-confidence samples at test time in addition to the adversarial training.
+Despite these advances, many methods rely on summary statistics (mean or median), direction (cosine similarity), or distance (Euclidean) of weight vectors.
+However, they often struggle to differentiate benign clients with non-IID data from malicious ones, leading to misclassification and suboptimal performance.
+The limited performance might be partially attributed to the inadequacy of Euclidean geometry for comparing neural network representations.
+In contrast, our approach uses Kernel CKA, a kernel-based metric, to measure the similarity between global and local models.
+This method effectively identifies compromised model parameters among clients with diverse data distributions, enhancing defense against model poisoning in FL and overcoming the shortcomings of previous techniques.
+2.3 Penultimate Layer Representation in FL
+The penultimate layer is a neural network model’s second-to-last layer (i.e., before the softmax layer).
 Wang et al.
-[22] discovered that PLRs are effective in distinguishing malicious models from benign ones; within benign clients, benign PLRs follow the same distribution, whereas malicious PLRs do not, across multiple datasets and neural network architectures.
+[30] discovered that PLRs are effective in distinguishing malicious models from benign ones; within benign clients, benign PLRs follow the same distribution, whereas malicious PLRs do not, across datasets and neural network architectures.
 Specifically, they demonstrated that distances among benign PLRs are smaller than those of benign and malicious PLRs.
-Their proposed method, FLARE, assigned a trust score to each client based on pairwise PLR discrepancies defined by Maximum Mean Discrepancy (MMD) among all model updates, allocating lower values to those farther from the benign distribution.
+Their proposed method, FLARE, assigned a trust score to each client based on pairwise PLR discrepancies defined by Maximum Mean Discrepancy among all model updates, allocating lower values to those farther from the benign distribution.
 The model updates are then scaled and averaged, weighted by each client’s trust score.
-However, FLARE continuously redirects the global model using server-owned raw data, increasing the risk of data leakage and jeopardizing data privacy.
-II-D Centered Kernel Alignment
+Meanwhile, FLARE continuously redirects the global model using server-owned raw data, increasing the risk of data leakage and jeopardizing data privacy.
+2.4 Centered Kernel Alignment
 Centered Kernel Alignment (CKA) [13] is a highly accurate similarity metric that measures how similar two differently initialized or trained neural networks are, producing a value between 0 (not similar) and 1 (identical).
-CKA can be calculated by:
-$CKA(X,Y)$ $=\frac{HSIC(X,Y)}{\sqrt{HSIC(X,X)HSIC(Y,Y)}}$
-$=\frac{{\lVert K(X^{T})K(Y)\rVert}^{2}}{{\lVert K(X^{T})K(X)\rVert}{\lVert K(Y^{T})K(Y)\rVert}}$
-where $HSIC$ is Hilbert-Schmidt Independence Criterion [10] (i.e., a not normalized version of CKA) and $K$ is the RBF kernel.
+CKA can be calculated by: $CKA(X,Y)=\frac{HSIC(X,Y)}{\sqrt{HSIC(X,X)HSIC(Y,Y)}}=\frac{{\lVert K(X^{T})K(Y)\rVert}^{2}}{{\lVert K(X^{T})K(X)\rVert}{\lVert K(Y^{T})K(Y)\rVert}}$ where $HSIC$ is Hilbert-Schmidt Independence Criterion [9] (i.e., a not normalized version of CKA) and $K$ is the RBF kernel.
 It constructs similarity kernel matrices in the weight space and compares these matrices to characterize the representation space, allowing comparisons between layers with varying widths and different initialization schemes.
-CKA includes several important properties for similarity measures of neural networks.
+CKA includes several vital properties for similarity measures of neural networks.
 It is non-invariant to invertible linear transformations, meaning similarity scores change if such transformations are applied.
@@
 However, CKA has yet to be leveraged as a defense mechanism against attacks involving models from malicious clients where the data are distributed across the clients in a non-IID manner.
-III Threat Model
-III-A Attacker’s Goal
+3 Threat Model
+3.1 Attackers’ Goals
 We consider two primary attack strategies in the context of FL: untargeted model poisoning and targeted backdoor attacks.
@@
 Targeted backdoor attacks, represented by Bhagoji et al.
-[3], aim to deceive the global model into misclassifying specific data samples the attacker chose.
+[3] and DBA [33], aim to deceive the global model into misclassifying specific data samples the attacker chose.
 The objective is to assign a different label chosen by the attacker while maintaining high accuracy for the remaining classes, making the attack more inconspicuous.
-III-B Attackers’ Capabilities
+3.2 Attackers’ Capability
   • An attacker on the client side can control multiple compromised clients.
   • An attacker has full control over at most $k<n/2$ clients.
-  • An attacker has knowledge about compromised clients’ data samples, current local model parameters, global model parameters of the previous round, and hyperparameters, such as learning rate, optimization method, loss function, batch size, and the number of local epochs.
+  • An attacker has knowledge about compromised clients’ data, such as current local model, previous global model, and hyperparameters.
   • An attacker has no knowledge or control over the server and honest clients.
-III-C Attack Strategy
-III-C1 Untargeted Model Poisoning Attacks
-These attacks based on [8] specifically aim to break Krum and Coordinate-wise Median (Coomed) aggregation rules, known to be byzantine failure tolerant.
+3.3 Attack Strategy
+3.3.1 Untargeted Model Poisoning Attacks [7]
+specifically aim to break Krum and Coordinate-wise Median (Coomed) aggregation rules, known to be byzantine failure tolerant.
 Their primary objective is to disrupt the model training process, thereby diminishing global test accuracy.
@@
 Specifically, we denote the attack against Krum as Untargeted-Krum and the attack against Coomed as Untargeted-Med.
-Untargeted-Krum [8] alters malicious parameters resembling benign ones to maximize the chances of being selected by Krum. Specifically, the optimization problem is as follows:
-$\max_{\lambda}\lambda$
-subject to $\text{subject to }w_{1}$ $=Krum(w_{1},...,w_{m},w_{m+1},...,w_{n})$
-$w_{1}$ $=w_{G}-\lambda s$
-$w_{i}$ $=w_{1},\text{ for }i=2,3,...,m$
-where $m$ is the number of attackers, $n$ is the total number of selected clients, $w_{G}$ is the global model from the previous iteration, and $s$ is the sign of the global model parameter with no attack.
-The upper bound of the $\lambda$ is solved as follows:
-$\lambda\leq$ $\frac{1}{(n-2m-1)\sqrt{d}}\cdot\min_{m+1\leq i\leq n}\sum_{l\in\Gamma_{w_{i}}^{n-m-2}}D(w_{l},w_{i})$
-$+\frac{1}{\sqrt{d}}\cdot\max_{m+1\leq i\leq n}D(w_{i},w_{G})$
-where $d$ is the number of parameters in the global model, $\Gamma$ is the set of $n-m-2$ benign local models having the smallest Euclidean distance to $w_{i}$, and $D$ is the Euclidean distance.
+Untargeted-Krum [7] alters malicious parameters resembling benign ones to maximize the chances of being selected by Krum.
+Specifically, the optimization problem is to find the maximum $\lambda,s.t.$ $w_{1}=Krum(w_{1},...,w_{m},w_{m+1},...,w_{n}),w_{1}=w_{G}-\lambda s,w_{i}=w_{1},$ for $i=2,3,...,m$ where $m$ is the number of attackers, $n$ is the total number of selected clients, $w_{G}$ is the previous global model, and $s$ is the sign of the global model parameter with no attack.
+The upper bound of the $\lambda$ can be solved as follows: $\lambda\leq\frac{1}{(n-2m-1)\sqrt{d}}\cdot\min_{m+1\leq i\leq n}\sum_{l\in\Gamma_{w_{i}}^{n-m-2}}D(w_{l},w_{i})+\frac{1}{\sqrt{d}}\cdot\max_{m+1\leq i\leq n}D(w_{i},w_{G})$ where $d$ is the number of parameters in the global model, $\Gamma$ is the set of $n-m-2$ benign local models having the smallest Euclidean distance to $w_{i}$, and $D$ is the Euclidean distance.
 $\lambda$ is halved until one of the compromised models is selected, or $\lambda$ is less than a threshold.
-Untargeted-Med [8] manipulates the model parameters based on their maximum and minimum so that chosen coordinate-wise median values direct toward an inverse direction.
-The attack starts with defining $w_{max,j}$ and $w_{min,j}$, which indicates the maximum and minimum of the $j$th local model parameters on the benign clients, respectively.
-$w_{max,j}$ $=\text{max}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$
-$w_{min,j}$ $=\text{min}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$
+Untargeted-Med [7] manipulates the model parameters based on the maximum and minimum so that chosen coordinate-wise median values direct toward an inverse direction.
+The attack starts with defining $w_{max,j}=\text{max}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$, which indicates the maximum, and $w_{min,j}=\text{min}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$ which indicates the minimum of the $j$th local model parameters on the benign clients, respectively.
 Also, to avoid sampled $m$ numbers being outliers, if $s_{j}=-1$, $m$ numbers in [$w_{max,j},b\cdot w_{max,j}$] (when $w_{max,j}>0$) or [$w_{max,j},b/w_{max,j}$] (when $w_{max,j}\leq 0$) are randomly sampled.
 Otherwise, $m$ numbers [$w_{min,j}/b,w_{min,j}$] (when $w_{min,j}>0$) or [$b\cdot w_{min,j},w_{min,j}$] (when $w_{min,j}\leq 0$) are randomly sampled.
-Since the attack does not depend on $b$, we set $b=2$ as the same as [8].
-III-C2 Targeted Backdoor Attack
-This attack is based on [3].
-Each malicious client owns one mislabeled image samples and trains the local model on it.
-We trained the model on backdoor tasks along with the main task(s); the training went on for both malicious and benign tasks to maintain benign accuracy such that side-job training (i.e., backdoor training) remains stealthy.
-Then, the attack boosted malicious clients’ updates to negate the combined effect of the benign client:
-$w_{i}^{t}=w_{G}^{t-1}+\alpha_{m}(w_{i}^{t-1}-w_{G}^{t-1})\text{ for }i=1,...,m$
-where $t$ is the current epoch and $\alpha_{m}$ is a boosting factor.
-In CIFAR10, for example, we manipulate the model parameters to misclassify an image of ’airplanes’ as ’bird’ whereas normally classifying the image of other classes like ’cat,’ ’horse,’ or ’truck.’
-IV FedCC: Robust Aggregation Against Poisoning Attacks
-Figure 2: An Overall Mechanism of our FedCC.
-IV-A Overview of FedCC
+We set $b=2$ as the same as the paper [7].
+3.3.2 Targeted Backdoor Attacks
+are based on [3].
+Each malicious client owns one sample of mislabeled images and trains the local model on it.
+We trained the model on backdoor tasks along with the main task(s); the training went on for both malicious and benign tasks to maintain benign accuracy such that backdoor training remains stealthy.
+Then we boosted malicious clients’ updates to negate the combined effect of the benign agent: $w_{i}^{t}=w_{G}^{t-1}+\alpha_{m}(w_{i}^{t-1}-w_{G}^{t-1})\text{ for }i=1,...,m$ where $t$ is the current epoch and $\alpha_{m}$ is a boosting factor.
+For DBA [33], we implant split backdoor triggers into the input so that the global model can be attacked in a distributed manner.
+For more detail, refer to [33].
+In CIFAR10, for example, we manipulate the model parameters to misclassify an image of ‘airplanes’ as ‘bird’ whereas normally classifying the image of other classes.
+4 FedCC: Robust Aggregation against Poisoning Attacks
+4.1 Overview of FedCC
 Figure 1 depicts an overview of FedCC.
-(1) A server initializes and advertises a global model to $n$ participating clients, selected out of total $k$ clients with a probability of fraction $C$.
+(1) A server initializes and advertises the global model to $n$ participating clients, selected out of total $k$ clients with a probability of fraction $C$.
 (2) The clients individually train their local models using their respective datasets and send the trained weights to the server for aggregation.
 Our proposed method intervenes at this moment.
-(3) The server extracts PLRs from the global model and each local model, and (4) compares RBF Kernel CKA values between them.
+(3) The server extracts PLRs from both the global model and each local model, and (4) compares RBF Kernel CKA values between them.
 (5) The CKA values are then clustered into two, potentially representing benign and malicious clients.
-(6) Since the server cannot assure whether the clients with higher similarities are malicious, it performs a coordinate-wise average of the weights from clients belonging to the cluster with a larger number of members, backed by an assumption that the number of malicious clients does not exceed $n/2-1$.
-(7) Finally, the server updates the global model for the current epoch and proceeds to the next iteration by distributing the updated global model to newly selected $n$ clients.
-IV-B Detailed Design
-When filtering malicious clients in the presence of non-IID data, a major challenge is the server’s lack of knowledge regarding the data distribution.
-Specifically, even if the data is benign, local models can exhibit significant angular or magnitude differences if the data is non-IID.
-We thus require a similarity metric independent of data distribution and not influenced by the distance or direction of model parameters.
+(6) The CKA similarity values are then normalized across all clients for every layer as $a$, but the second last layer’s CKA similarity values are normalized within the larger cluster as $b$, backed by the assumption that the number of malicious clients does not exceed $n/2-1$.
+(7) Finally, the server aggregates each layer of local models weighted by $a$, but the PLR weighted by $b$, to update a global model for the current epoch and proceeds to the next epoch by distributing the updated global model to the newly selected $n$ clients.
+Figure 1: An Overview of FedCC.
+$k$, $n$, $a$, and $b$ indicate the kth layer, the number of participating clients, normalized similarity across all layers (i.e., across_cka), and normalized similarity within the cluster (i.e., within_cka), respectively.
+4.2 Detailed Design
+A major challenge when filtering malicious clients in the presence of non-IID data is the server’s lack of knowledge regarding the data distribution.
+Even if the data is benign, local learning models can exhibit significant angular or magnitude differences if they are non-IID.
+Thus, we require a similarity metric independent of data distribution and not influenced by the distance or direction of model parameters.
+4.2.1 Penultimate Layer
+is the output of the second-to-last layer before the softmax layer in CNN.
+In federated learning, where the server cannot access local training data, it receives trained model parameters (weights) from selected clients.
+Therefore, the collected weights are the only available information for determining whether a client is malicious.
+PLR takes the end result of the convolution and pooling process that will reach a classification decision.
+Thus, we assumed that computed weights of the second last layer might contain the essential information that strongly influences classification results.
 Figure 2: Cluster Distance of Each Layer
-TABLE I: Comparison of Performance with Various Similarity Metrics
+To justify the utilization of PLRs, we measure the distance between two clusters of clients (benign and malicious) using dendrograms with a single linkage metric and correlation distance method.
+Figure 2 presents cluster distances of each layer under different attack scenarios, such as untargeted-mKrum attacks in IID or non-IID settings.
+Observably, the cluster distance is the most considerable in PLR to other layers, indicating the highest degree of differences.
+Based on this observation, we conclude that PLR contains the most distinguishable and indicative information.
+Therefore, we utilize the weights of the penultimate layer to compare their similarity to the global model.
+4.2.2 Kernel CKA[13]
+is a similarity metric to assess the similarity or dissimilarity between two neural networks.
+It compares representations at different layers or models trained in different ways.
+The sophisticated similarity comparison ability of Kernel CKA makes it a suitable choice for our purpose.
+To justify the effectiveness of Kernel CKA, we compared its performance with other similarity metrics such as linear CKA, MMD (not normalized CKA), cosine similarity (angle), and Euclidean similarity (distance).
+We evaluated the test accuracy of FedCC using the CIFAR10 dataset under two untargeted attacks and one targeted attack, separately in IID and non-IID settings.
+The results are summarized in Table 1.
+In the ’targeted’ column, the values represent the combination of test accuracy and backdoor confidence, with higher test accuracy and lower backdoor confidence indicating better performance.
+We observed that Kernel CKA consistently yields the highest performance across experiments; therefore, we decided to utilize Kernel CKA to compare the similarities of PLRs.
+Table 1: Comparison of Performance with Various Similarity Metrics
 Method Fang-Med Fang-mKrum Targeted
@@
 Euclidean 69.06 27.82 68.54 41.57 69.17/0.0221 52.20/0.0015
-Penultimate Layer is the output of the second-to-last layer before the softmax layer in CNN.
-In federated learning, where the server cannot access local training data, it receives trained model parameters (weights) from selected clients.
-Therefore, the collected weights are the only available information for determining whether a client is malicious.
-PLR takes the end result of the convolution and pooling process that will reach a classification decision.
-Thus, we assumed that computed weights of the second last layer might contain the essential information that strongly influences classification results.
-To justify the utilization of PLRs, we measure the distance between two clusters of clients (benign and malicious) using dendrograms with a single linkage metric and correlation distance method.
-Figure 3 presents cluster distances of each layer under different attack scenarios, such as untargeted-mKrum attacks in IID or non-IID settings.
-Observably, the cluster distance is the most considerable in PLR to other layers, indicating the highest degree of differences.
-Based on this observation, we conclude that PLR contains the most distinguishable and indicative information.
-Therefore, we utilize the weights of the penultimate layer to compare their similarity to the global model.
-Kernel CKA[13] is a similarity metric to assess the similarity between two neural networks.
-It compares representations at different layers or models trained in various ways.
-The sophisticated similarity comparison ability of Kernel CKA makes it a suitable choice for our purpose.
-To justify the use of Kernel CKA, we compared its performance with other similarity metrics such as linear CKA, MMD (not normalized CKA), cosine (angle), and Euclidean similarity (distance).
-We evaluated the test accuracy of FedCC using the CIFAR10 dataset under two untargeted attacks and one targeted attack, separately in IID and non-IID settings.
-The results are summarized in Table I.
-In the ’targeted’ column, the values represent the combination of test accuracy and backdoor confidence, with higher test accuracy and lower backdoor confidence indicating better performance.
-We observed that Kernel CKA consistently yielded the highest performance across multiple experiments.
-Therefore, we decided to utilize Kernel CKA to compare PLR similarities.
-Algorithm 1 FedCC
-Input global_w, n local_w $w_{1}$, …, $w_{n}$
-      Output agg_w
-1:for $i<n$ do $\triangleright$ Get PLRs of local and global models
-2:     local_plrs[$i$] $\leftarrow$ local_ws[second_last_layer]
-3:end for
-4:glob_plr $\leftarrow$ global_w[second_last_layer]
-5:for $i<n$ do $\triangleright$ Apply kernel CKA to the plrs
-6:     cka[$i$] $\leftarrow$ cka(glob_plr, local_plrs[$i$])
-7:end $\triangleright$ Apply Kmeans clustering algorithm
-8:kmeans = kmeans(n_cluster=2, cka[$i$])
-9:labels = kmeans.labels_
-10:count = counter(labels)
-11:suspect $\leftarrow$ 1
-12:if count[00] $\leq$ count[$1$then
-13:     suspect $\leftarrow$ 0
-14:end $\triangleright$ Define suspect based on the size of cluster
-15:suspects = where(label == suspect)
-16:for $i$ in range($n$do
-17:     for $s$ in suspects do $\triangleright$ Set scale to 0 for the suspects
-18:         scale[$s$] $\leftarrow 0$
-19:     end for
-20:     if scale[$i$] != 0 then
-21:         scale[$i$] $\leftarrow 1$
-22:         selected_param.append(scale[$i$] * local_ws)
-23:     end if
-24:end $\triangleright$ Average the selected parameters
-25:agg_weights = mean(scale*selected_param)
-26:return agg_weights
-FedCC is an aggregation method that combines selected weights without scaling, based on the measurement of Kernel CKA similarities between the global and local models’ PLRs.
-A pseudo algorithm is described in Algorithm 1.
+4.2.3 FedCC
+is an aggregation method that combines local clients’ model parameters, based on the Kernel CKA similarities between the global and local models’ PLRs.
+An overall algorithm is described in Appendix 0.B.
 A server first extracts PLRs of the global and local models and calculates the kernel CKA between them.
@@
 Any clustering method that produces only two clusters can be used.
-Considering the assumption that the number of attackers is less than half of the total number of clients ($n/2$), we designate the clients in the cluster with fewer members as suspects.
-The indices of these suspicious clients are stored in a list called ‘suspects.’ We proceed by selectively averaging the weights of clients whose index does not belong to the ‘suspects’ list, excluding the weights of the suspicious clients from the aggregation.
-Note that the weights are averaged without scaling, as scaling could impede the model from convergence even when the loss reaches optima.
-We named our proposed method FedCC.
-Compared to other methods such as Krum, Bulyan, and FLARE, which either require prior knowledge of the number of attackers or allow the server to have access to raw data, FedCC does not rely on any information about the clients or any instances of data.
-This ensures strong privacy preservation within the federated learning framework.
-Furthermore, FedCC effectively distinguishes between malicious and benign clients, even in practical but challenging scenarios where benign clients have non-IID data.
-V Experiments
-V-A Dataset
-We use three benchmark vision datasets: Fashion-MNIST (fMNIST) [24], CIFAR10, and CIFAR100 [14].
-fMNIST consists of 60,000 training and 10,000 testing samples of 28$\times$28 grayscale images and associating labels of 10 classes.
-The labels include T-shirt/top, Trouser, Pullover, Dress, Coat, Sandal, Shirt, Sneaker, Bag, and Ankle boot.
-CIFAR10 consists of 50,000 training and 10,000 testing examples of 32$\times$32 color images and associating labels of 10 classes.
-The classes include airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck.
-CIFAR100 is the same as CIFAR10 except for the categorization.
-It is labeled as 100 classes, each containing 600 images.
-Thus, there are 500 training and 100 testing images per class.
-Each class (i.e., ‘fine’ label) belongs to one of the 20 superclasses.
-V-B Non-IID Simulation
-Non-IID is a more feasible assumption considering the diverse and massive nature of data in practice.
-To standardize non-IID settings, we use Dirichlet distribution, which is well-known for its capability of representing real-world data [15].
-A concentration parameter $\alpha$ controls the degree of non-IID; the smaller $\alpha$ is, the more likely the clients hold examples from only one randomly chosen class.
-V-C Model Architecture
+Considering the assumption that the number of attackers is less than half of the total number of clients ($n/2$), we designate the clients in the larger cluster as ‘candidates.’
+Since the earlier layers in CNNs learn global features and the latter layers embed local features [15, 25, 23], we used two differently normalized CKA similarity values: one normalized within the candidates (i.e., within_cka, denoted as $b$ in Figure 1) and one normalized across all the clients (i.e., across_cka, denoted as $a$ in Figure 1).
+For the PLR (second-last layer), we apply within_cka to the clients belonging to the larger cluster, whereas for all other layers, we apply across_cka to weighted average model parameters.
+By doing so, we can ensure that the end magnitude of the aggregated parameters stays similar to before the manipulations, leveraging the fact that the sum of the normalized values is one.
+Moreover, since our method evaluates clients in every global round and does not exclude any local models, it is also effective for adaptive attacks like DBA.
+5 Experiments
+5.1 Dataset and Model Architecture
+We use three benchmark vision datasets: Fashion-MNIST (fMNIST) [32], CIFAR10 and CIFAR100 [14].
+Details about datasets can be found in Appendix 0.A.
 Considering that the end devices are normally incapable of handling heavy computation due to resource or communication constraints, we used lightweight CNN architectures for the experiments.
@@
 For CIFAR10, we utilized a 5-layer-CNN with three convolutions, each followed by a max-pool layer and two FC layers, applying dropout with a probability of 0.5 before the first FC layer.
-Finally, for CIFAR100, we adopted the LeNet architecture.
-Table VI, Table VII, and Table VIII in Appendix summarize model architectures for each benchmark dataset.
-V-D Setup
-We implement two untargeted attacks in [8] and one targeted attack [3], defense baselines, including Krum [4], Coomed [26], Bulyan [11], and FLARE [22], and FedCC.
-We have ten clients with a client participation fraction of 1.0 unless stated otherwise.
-The number of attackers is 30% of the total clients (three) in untargeted and 10% (one) in targeted attacks as in [22].
-We also assumed all clients have the same NN architecture.
-Under the targeted backdoor attack, benign clients go through three, whereas compromised clients undergo six local epochs.
-We train each client’s local model using an Adam optimizer with a learning rate of 0.001, following the experiment in [22].
-All results are averaged after running each experiment three times.
-TABLE II: Accuracy of no-attack scenario, boldface and underlining refer to the highest and the second highest accuracy, respectively.
-Scenarios Dataset FedAvg Krum Coomed Bulyan FLARE FedCC
-No Attack non-IID fMNIST 69.68 43.10 76.12 68.78 42.82 71.92
-CIFAR10 53.06 16.16 45.55 46.67 27.33 50.80
-CIFAR100 17.22 7.27 12.82 15.21 0.99 17.62
-No Attack IID fMNIST 89.89 83.81 88.91 89.49 76.86 89.71
-CIFAR10 69.17 56.56 67.34 67.81 46.84 69.03
-CIFAR100 21.49 7.73 15.56 20.55 7.59 22.05
-V-E Evaluation Metrics
-We assessed the performance of our defense mechanism using two metrics: backdoor confidence (confidence) and global model test accuracy (accuracy).
-The confidence indicates the likelihood of misclassification, with values ranging from 0 to 1 (lower values indicating better defense performance).
-Accuracy measures the overall test accuracy of the global model when various defense methods are applied when attacks exist, with values ranging from 0 to 100 (higher values indicating better performance).
-We reported both metrics in the presence of backdoor attacks while focusing solely on accuracy for untargeted attacks—as the attacker’s goal is to deteriorate the overall accuracy.
-VI Results and Discussion
-In Table II, which represents the baseline experiments without any attacks, we can observe the accuracy of different methods in both non-IID and IID settings.
-FedCC achieves accuracy comparable to that of the best-performing method, FedAvg.
-The slightly higher accuracy of FedAvg than FedCC may be attributed to its wholesome averaging, which leads to a more generalized model when there is no attack.
-Nonetheless, FedCC surpasses all other baseline methods with CIFAR100, which has more classes, making it harder to achieve better accuracy in FL[7].
-TABLE III: Accuracy under untargeted and targeted attacks in Non-IID (top) and IID setting (bottom)
-Scenarios Dataset FedAvg Krum Coomed Bulyan FLARE FedCC
-Fang-Krum non-IID fMNIST 64.63 16.51 57.12 13.39 49.54 66.27
-CIFAR10 33.69 15.38 35.70 19.23 17.03 37.40
-CIFAR100 2.27 1.00 4.85 0.98 7.46 14.51
-Fang-Med non-IID fMNIST 16.32 49.33 68.90 64.12 52.25 72.76
-CIFAR10 10.02 25.06 40.23 32.47 14.59 42.37
-CIFAR100 1.00 6.24 10.27 6.91 1.00 16.12
-Targeted non-IID fMNIST 79.07 43.46 71.25 73.58 60.57 84.54
-CIFAR10 30.28 16.05 28.87 25.77 10.52 36.20
-CIFAR100 9.84 7.24 11.02 12.01 1.04 13.81
-Fang-Krum IID fMNIST 75.55 31.66 87.62 50.30 79.16 89.23
-CIFAR10 49.67 40.86 57.40 12.67 25.77 67.12
-CIFAR100 13.72 1.04 6.17 1.59 7.49 18.47
-Fang-Med IID fMNIST 20.86 85.33 86.70 87.45 71.08 88.44
-CIFAR10 9.51 54.28 59.20 57.69 49.82 64.74
-CIFAR100 0.87 12.27 14.43 12.56 5.83 17.83
-Targeted IID fMNIST 89.32 85.52 88.45 89.35 67.65 90.16
-CIFAR10 52.20 45.37 49.27 55.77 10.09 51.57
-CIFAR100 12.73 8.37 15.91 16.59 1.71 18.08
-VI-A Non-IID Data Environment
-Following the aforementioned non-IID simulation, we set $\alpha$ to 0.2 to create a disjoint client training data and distribute each set of data to each client.
-VI-A1 Untargeted Model Poisoning Attacks
-The upper section of Table III provides the global model’s accuracy across three datasets under two untargeted (Fang-Krum and Fang-Med) and Targeted attacks in a non-IID environment.
-One notable finding is that robust aggregation algorithms often yield lower accuracy compared to simple averaging (no defense).
-This discrepancy arises from the imperfect identification of malicious clients, leading to the erroneous aggregation of their weights along with benign clients.
-When applying Krum and Bulyan against Fang-Krum, the accuracy drops by a substantial 59.30% and 65.68%, respectively, compared to FedAvg.
-This decline can be attributed to Fang-Krum’s intentional design to degrade accuracy against Krum.
-Additionally, Coomed struggles to effectively mitigate this attack as the median values of weights often fail to represent benign clients, particularly in non-IID settings, accurately.
-Furthermore, the relatively low accuracy of FLARE is due to weight scaling, even when the model already exhibits satisfactory performance.
-In contrast, FedCC, which selectively averages weights, stands out by neither boosting nor minimizing weights, resulting in higher performance than other methods.
-Notably, FedCC achieves the highest accuracy (66.27%, 37.40%, and 14.51% for fMNIST, CIFAR10, and CIFAR100, respectively), with a significant leap in experiments with CIFAR100.
-Turning our attention to the Fang-Med attack, the accuracy of FedAvg experiences a substantial drop due to the attack’s tendency to deviate from median values, creating outliers.
-This behavior explains the relatively high accuracy when Coomed and Bulyan are applied because coordinate-wise median and trimmed mean methods disregard outliers.
-Meanwhile, FLARE performs poorly for a reason similar to the one mentioned above.
-For both attacks, we can observe that Krum, which selects a single local model as the global model, is less robust than Coomed, which aggregates multiple local models.
-Similarly, Bulyan, which takes averages of multiple chosen clients, exhibits less robustness than Coomed, particularly in non-IID data environments.
-Notably, FedCC achieves the highest accuracy (72.76%, 42.37%, and 16.12% for each dataset) without sharing raw data with the server or other clients with non-IID data.
-This exceptional performance of FedCC can be attributed to two factors: 1) highly distinguishable information within PLRs, and 2) higher CKA similarity between benign clients than the one between benign and malicious clients.
-VI-A2 Targeted Backdoor Attacks
-Figure 4: Confidence of Backdoor Task in a Non-IID Setting (top) and an IID setting (bottom)
-The fourth row of Table III summarizes accuracy under targeted backdoor attacks in a non-IID setting.
-To further illustrate the impact of these attacks, we include a visual representation of the backdoor confidence in the first row of Figure 4.
-It becomes evident that Krum, which selects a single client’s weights as the global model, exhibits diminished robustness compared to other methods that aggregate the weights of multiple clients; coordinate-wise median (Coomed) and trimmed mean with multi-Krum (Bulyan) demonstrate relatively higher test accuracy.
-Conversely, FLARE decreases the main task test accuracy for similar reasons observed in the experiments mentioned above.
-Nonetheless, FedCC achieves the highest test accuracy.
-This outcome underscores the superiority of filtering malicious clients through CKA similarity comparison, surpassing the effectiveness of first-order statistical methods like mean or median.
-Notably, Figure 4 illustrates how FedCC significantly nullifies backdoor confidence.
-Compared to oscillating values of other methods, the stable low confidence values further emphasize FedCC’s resilience and independence from the specific data distribution across clients.
-This advantage stems from its effective utilization of CKA, enabling the calculation of similarity values even when models are trained on different datasets with varying data distributions.
-Whereas it is true that other baselines, such as Bulyan, also reduce confidence, we argue that FedCC not only preserves but potentially increases accuracy by precisely discriminating against malicious clients.
-VI-B IID Data Environment
+For CIFAR100, we adopted LeNet [16].
+5.2 Non-IID Simulation
+Non-IID is a more feasible assumption considering the diverse and massive nature of data in practice.
+To standardize non-IID settings, we use Dirichlet distribution, which is widely used and well-known for its capability of representing real-world data [18].
+A concentration parameter $\alpha$ controls the degree of non-IID; the smaller $\alpha$ is, the more likely the clients hold examples from only one randomly chosen class.
+We set the $\alpha$ to 0.2 to represent non-IID, based on [11, 38].
+5.3 Experimental Setup and Baselines
+We implement two untargeted attacks from [7] and two targeted attacks from [3] and [33], along with defense baselines including Krum, multi-Krum [4], Coomed [34], Bulyan [10], FLTrust [6], FLARE [30], and FedCC.
+We use ten clients with a 1.0 client participation fraction unless stated otherwise.
+As in [30], we have three adversaries for each untargeted attack and one for the targeted attack.
+In the DBA, we randomly select four malicious clients and implant five distributed triggers per batch.
+To compensate for delayed attack effects, we use FedAvg until epoch 30, after which each defense method is applied.
+Note that, for a fair comparison, FLTrust is computed using its previous global model without auxiliary data.
+Benign clients undergo three local epochs, while compromised clients undergo six, with training done using the Adam optimizer (learning rate = 0.001).
+5.4 Evaluation Metrics
+We evaluated our defense mechanism using two metrics: backdoor confidence (confidence) and global model test accuracy (accuracy).
+The confidence metric measures misclassification likelihood, ranging from 0 to 1 (lower values are better).
+Accuracy reflects the global model’s overall test performance under defense methods during attacks, ranging from 0 to 100 (higher values are better).
+Both metrics were reported under backdoor attacks, with a focus on accuracy for untargeted attacks, as the attacker’s goal is to reduce overall accuracy.
+6 Results and Discussion
+In this section, we report the performance of our FedCC against two untargeted and two targeted attacks in both IID and non-IID environments.
+We additionally demonstrate the defense effect against a distributed adaptive attack, DBA [33].
+Unlike centralized backdoor attacks, the DBA splits a trigger pattern across multiple adversaries, evading the detection by anomaly detection.
+We show the result of CIFAR10, with various defense methods applied.
+6.1 Non-IID Data Environment
+Non-IID is a more feasible assumption considering the diverse and massive nature of data in practice.
+To standardize non-IID settings, we employ Dirichlet distribution that 0 indicates the most heterogeneity, and 100 mimics homogeneity [20].
+A concentration parameter $\alpha$ controls the degree of non-IID; the smaller $\alpha$ is, the more likely the clients hold examples from only one randomly chosen class.
+Since $\alpha$ being 0.2 represents a highly non-IID scenario based on [11, 38], we set it as such, following [38].
+Table 2: Test Accuracy under untargeted attacks in Non-IID setting.
+Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
+Fang -Krum non-IID fM 64.63 16.51 45.91 57.12 13.39 49.37 49.54 66.27
+C10 33.69 15.38 20.09 35.70 19.23 36.63 17.03 37.40
+C100 2.27 1.00 5.34 4.85 0.98 11.62 7.46 14.51
+Fang -Med non-IID fM 16.32 49.33 66.30 68.90 64.12 12.38 52.25 72.76
+C10 10.02 25.06 40.52 40.23 32.47 12.39 14.59 42.37
+C100 1.00 6.24 13.97 10.27 6.91 1.00 1.00 16.12
+6.1.1 Untargeted Model Poisoning Attacks in non-IID setting
+Table 2 provides the global model’s accuracy across three datasets under two untargeted (Fang-Krum and Fang-Med) in a non-IID environment.
+A notable finding is that robust aggregation algorithms often yield lower accuracy than simple averaging (fedAvg).
+This discrepancy arises from the imperfect identification of malicious clients, leading to the erroneous aggregation of their weights along with those of benign clients.
+Under Fang-Krum, Krum, Bulyan, and multi-Krum experience significant accuracy drops (16.51%, 13.39%, and 45.91%, respectively, compared to FedAvg’s 64.63%), reflecting the attack’s design to degrade Krum-based defenses.
+Coomed also struggles, as median weights fail to represent benign clients effectively in non-IID settings.
+FLTrust mitigates Fang-Krum but performs better on CIFAR10 than fMNIST, likely due to fMNIST’s simpler gradients and lower variance, which challenge its ability to distinguish malicious updates.
+FLARE’s lower accuracy stems from its approach of scaling the entire model weights uniformly, which ignores knowledge contributed by individual clients.
+In contrast, FedCC, which selectively averages weights layer-wise, achieves superior performance by boosting or minimizing weights based on their alignment with benign updates.
+FedCC demonstrates the highest accuracy (66.27%, 37.40%, and 14.51% for fMNIST, CIFAR10, and CIFAR100, respectively) and shows significant improvement in CIFAR100 experiments.
+Turning to the Fang-Med attack, FedAvg experiences a significant drop in accuracy due to the attack’s creation of outliers by deviating from median values.
+Consequently, Coomed and Bulyan perform relatively well since their coordinate-wise median and trimmed mean methods disregard outliers effectively.
+FLARE continues to perform poorly for a similar reason mentioned earlier.
+Between the Krum-based methods, multi-Krum and Coomed are more robust than Krum, as they aggregate multiple local models, providing better resilience to outliers.
+Similarly, Bulyan’s averaging of multiple clients enhances robustness but falls short of Coomed, particularly in non-IID environments.
+Finally, FedCC achieves the highest accuracy (72.76%, 42.37%, and 16.12% for each dataset) without sharing raw data, showcasing its exceptional performance.
+This success is attributed to two factors: (1) highly distinguishable information within PLRs and (2) higher CKA similarity between benign clients than between benign and malicious clients.
+6.1.2 Targeted Backdoor Attacks
+Table 3: Test Accuracy under targeted attacks in Non-IID setting.
+Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
+Target non-IID fM 75.65 45.27 65.97 71.70 57.96 61.82 64.31 75.66
+C10 36.16 14.98 30.72 48.97 40.11 44.06 10.18 51.56
+C100 4.46 6.18 6.90 12.04 11.16 12.95 1.14 15.26
+DBA C10 38.56 24.94 7.09 44.45 34.19 51.49 38.73 52.28
+Figure 3: Confidence of Backdoor Task for targeted attacks in Non-IID settings.
+Table 3 summarizes test accuracy under targeted attacks in a non-IID setting.
+To further illustrate the impact, Figure 3 provides a visual representation of backdoor confidence.
+It is evident that Krum, which selects a single client’s weights as the global model, shows reduced robustness compared to methods that aggregate weights from multiple clients.
+In contrast, Coomed, multi-Krum, and Bulyan exhibit higher test accuracy.
+FLARE reduces main task accuracy for reasons similar to those observed in previous experiments.
+FedCC achieves the highest test accuracy.
+Since targeted attacks aim to misclassify a specific target class while maintaining the main task accuracy, the best performance of FedCC is attributed to leveraging knowledge from earlier stages of training.
+This result highlights the superiority of using sophisticated weighted knowledge and filtering malicious clients via CKA similarity, which outperforms first-order statistical methods like mean or median.
+A similar trend is observed with the DBA attack, except for multi-Krum; the lowest accuracy occurs because all four malicious clients, each with a distributed trigger, are treated as outliers and excluded from aggregation.
+Notably, Figure 3 demonstrates that FedCC significantly reduces backdoor confidence.
+Note that since the backdoor confidence of DBA fluctuates a lot due to its distributed nature, we omit it for brevity.
+Unlike other methods, such as FLTrust or multi-Krum, which exhibit fluctuating confidence values, FedCC maintains consistently low confidence, emphasizing its resilience and independence from client-specific data distributions.
+This advantage stems from the effective utilization of CKA, enabling similarity calculations even when models are trained on different datasets with varying distributions.
+While other baselines, like Bulyan or FLTrust, also reduce confidence, we argue that FedCC not only preserves but potentially increases accuracy by precisely distinguishing malicious clients.
+6.2 IID Data Environment
 In an IID setting, we evenly divide the training dataset among all clients so that each client’s data distribution is identical and of the same size.
-VI-B1 Untargeted Model Poisoning Attacks
-The bottom section of Table III (fifth to seventh rows) provides valuable insights into the test accuracy of a global model trained on three datasets under untargeted attacks in IID settings.
-Although the trends observed in non-IID experiments appear similar, the accuracy values are slightly higher in this case due to the identical nature of the data.
-Under the Fang-Krum attack, we can observe relatively low accuracy values for Krum and Bulyan, as the attack explicitly targets Krum and its variations.
-However, when Coomed is applied, the accuracy increases from 75.55% (FedAvg) to 79.16% (Coomed), indicating that coordinate-wise median values closely align with the weights of the global model with minimal deviation.
-FLARE’s low accuracy is due to the continuous scaling of weights even when the model is already optimized.
-In contrast, FedCC does not scale weights but focuses solely on optimizing the model to the minimum by averaging chosen clients.
-As a result, FedCC achieves the highest accuracy (89.23%, 67.12%, and 14.51% for each dataset) by accurately identifying and filtering out compromised clients.
+Table 4: Test Accuracy under untargeted attacks in IID setting.
+Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
+Fang -Krum IID fM 75.55 31.66 87.78 87.62 50.30 89.53 79.16 89.57
+C10 49.67 40.86 63.42 57.40 12.67 68.25 25.77 69.84
+C100 13.72 1.04 7.64 6.17 1.59 17.14 7.49 18.47
+Fang -Med IID fM 20.86 85.33 89.53 86.70 87.45 21.36 71.08 89.66
+C10 9.51 54.28 69.68 59.20 57.69 9.92 49.82 70.52
+C100 0.87 12.27 16.52 14.43 12.56 1.16 5.83 17.83
+6.2.1 Untargeted Model Poisoning Attacks
+Table 4 shows the test accuracy of the global model trained on three datasets under untargeted attacks in IID settings.
+While trends in these experiments are similar to those in the non-IID setting, the accuracy values are slightly higher due to the more consistent data distribution across and updates from clients in the IID case.
+A notable deviation from the non-IID experiments is when Coomed is applied: the accuracy increases from 75.55% (FedAvg) to 87.62% (Coomed).
+This indicates that the coordinate-wise median values closely align with the global model’s weights with minimal deviation and represent benign models well, as the adversary’s impact is limited due to the majority of benign clients.
+FedCC achieves the highest accuracy across all methods (89.57%, 69.84%, and 18.47% for each dataset) due to its effective CKA similarity measurement and sophisticated layer-wise aggregation strategy.
 Under the Fang-Med attack, the accuracy of FedAvg experiences a significant drop (89.89% to 20.86%) due to the peculiarity of outliers when the clients have IID data.
-Notably, Krum, Coomed, and Bulyan demonstrate similar accuracy values, indicating that coordinate-wise median values indeed represent benign clients’ parameters.
-Meanwhile, FedCC consistently achieves the highest accuracy, with values of 88.44%, 64.74%, and 17.83% for each dataset.
-Notably, the remarkable growth observed in experiments with CIFAR100, with a 66% and 58.7% increment under Fang-Krum and Fang-Med attacks, respectively, surpasses the second-highest improvement.
-These results strongly validate the effectiveness of FedCC in mitigating untargeted model poisoning attacks in a simple IID setting.
-VI-B2 Targeted Backdoor Attacks
-The last row of Table III and the bottom three figures of Figure 4 are the test accuracy and backdoor confidence under targeted backdoor attacks in an IID setting.
-In targeted backdoor attacks, the global model is covertly trained to misclassify a specific attacker-chosen class to an attacker-chosen label while leaving the remaining classes intact.
-As a result, the test accuracy has remained relatively high across the different methods.
-However, a critical observation is the remarkable confidence reduction achieved by FedCC.
-The backdoor confidence with FedCC approaches zero, effectively neutralizing the backdoor task.
-In comparison, the attack confidence of FedAvg and Bulyan, despite exhibiting higher test accuracy for CIFAR10, remains higher than that of FedCC.
-Note that the confidence of the other methods fluctuates by a large margin.
-Specifically, FLARE’s backdoor confidence oscillates even though it occasionally records lower confidence than ours (e.g., CIFAR10).
-This means that its performance is less stable than that of FedCC, which consistently minimizes backdoor confidence.
-It is crucial to emphasize that reducing attack confidence is as significant as maintaining the main test accuracy since the attacker’s objective is to train the model on backdoor tasks stealthily, but the defender’s goal is to nullify them.
-Hence, our proposed method achieves the highest accuracy for fMNIST and CIFAR100 and comparable accuracy for CIFAR10 while effectively nullifying the backdoor confidence.
-This outcome underscores the effectiveness and robustness of our approach in mitigating targeted backdoor attacks.
-VII Performance in Varying FL Settings
-We additionally measure the effectiveness of FedCC in various non-IID settings, including different numbers of malicious clients, fractions of participation, and numbers of local epochs.
-Since Coomed was as effective as ours, and FLARE also used PLRs, we compare the three methods to FedCC, including FedAvg.
-VII-A Impact of the Number of Malicious Clients
-Figure 5: Accuracy with different numbers of malicious clients
-We assess the impact of different numbers of malicious clients under untargeted attacks, specifically Fang-Krum and Fang-Med.
-In this set of experiments, we fix the total number of clients and participation fraction at 10 and 1, respectively.
-To adhere to the assumption that the number of attackers is less than half of the participating clients, we investigate the test accuracy considering scenarios involving up to four malicious clients.
+Similar to the result of Fang-Krum, Krum, Coomed, and Bulyan demonstrate similar accuracy values, indicating that coordinate-wise median values indeed represent benign clients’ parameters.
+FLTrust, however, is less effective in mitigating the Fang-Med attack than Fang-Krum, even in the IID setting; it then means that the performance difference is not due to data distribution but rather its inherent vulnerability.
+In Fang-Krum, malicious updates are deliberately far from the global model, allowing FLTrust to identify discrepancies easily.
+In contrast, the malicious updates in Fang-Med are subtler and only slightly deviate from the global model or median, making it harder for FLTrust to distinguish them from benign updates.
+Meanwhile, FedCC consistently achieves the highest accuracy (89.66%, 70.52%, and 17.83% for each dataset), demonstrating its capability to mitigate attacks involving more nuanced changes to model updates without extreme outliers.
+6.2.2 Targeted Backdoor Attacks
+Table 5: Test Accuracy under targeted attacks in IID setting.
+Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
+Target IID fM 88.27 86.63 87.03 89.41 89.45 89.59 75.29 90.01
+C10 64.68 57.69 71.19 69.85 68.76 68.61 11.09 71.64
+C100 13.83 5.72 17.57 13.24 15.11 17.08 1.03 18.61
+DBA C10 10.00 35.58 51.40 10.00 16.16 56.69 10.00 58.04
+Figure 4: Confidence of Backdoor Task for targeted attacks in IID settings.
+Table 5 and Figure 4 present the test accuracy and backdoor confidence under targeted backdoor attacks in IID settings.
+Similar to the untargeted attack scenario, the general test accuracy in IID settings is higher than in non-IID settings.
+Reducing backdoor confidence is as critical as maintaining main task accuracy, as attackers aim to stealthily embed backdoor tasks.
+A key observation is FedCC’s remarkable ability to reduce backdoor confidence to near zero while preserving high main task accuracy.
+In contrast, the fluctuating backdoor confidence observed in Multi-Krum and Flare is largely due to the distributed trigger, which confuses their defense mechanisms and hampers their ability to filter out outliers effectively.
+Other methods, such as Multi-Krum, Flare, and FLTrust, show significant variability in backdoor confidence, further highlighting FedCC’s superiority.
+Notably, FedCC achieved the highest main task accuracy across both targeted backdoor attacks and distributed backdoor attacks (DBA), consistently neutralizing the backdoor task.
+These results underscore FedCC’s ability to mitigate targeted backdoor attacks through precise and efficient integration of prior knowledge to preserve main task performance while zeroing out the attack confidence.
+6.3 Performance in Varying FL Settings
+We additionally measured the effectiveness of FedCC in various non-IID settings, including different numbers of malicious clients, fractions of participation, and numbers of local epochs.
+Since Coomed was as effective as ours, and FLARE also used PLRs, we compared three methods to FedCC, including FedAvg.
+6.3.1 Impact of the Number of Malicious Clients
+Figure 5: Test Accuracy with different numbers of malicious clients
+We assessed the impact of different numbers of malicious clients under untargeted attacks, specifically Untargeted-Krum and Untargeted-Med.
+In this set of experiments, we kept the total number of clients and participation fraction fixed at 10 and 1, respectively.
+To adhere to the assumption that the number of attackers is less than half of the participating clients, we investigated the test accuracy considering scenarios involving up to four malicious clients.
 Figure 5 illustrates the accuracy in the given experimental settings, and it is evident that FedCC outperforms the other defense methods.
 We observe a general trend where the accuracy of defense methods decreases as the number of attackers increases.
-Notably, Coomed experiences the most significant drop in accuracy when transitioning from one to four malicious clients.
-This can be attributed to the fact that as the participation of malicious clients increases, there is a higher likelihood that the median values will be influenced by the malicious contributions.
+Notably, Coomed experiences the most significant drop in accuracy when transitioning from one malicious client to four malicious clients.
+This can be attributed to the fact that as the participation of malicious clients increases, there is a higher likelihood of the median values being influenced by their malicious contributions.
 In contrast, FedCC does not rely solely on geometric measures such as angles or distances but instead leverages hidden correlations between networks to identify and filter out malicious clients.
@@
 These findings underscore the robustness and effectiveness of FedCC in defending against untargeted attacks, as it consistently outperforms other defenses across varying numbers of malicious clients.
-VII-B Impact of Fraction
-TABLE IV: Accuracy under untargeted attacks with different fractions of clients being selected. fM refers to fMNIST, C10 refers to CIFAR10, Med refers to Coordinate wise median.
-Fang-Krum Fang-Med
+6.3.2 Impact of Fraction
+Table 6: Test Accuracy under untargeted attacks with different fractions of clients being selected.
+Untargeted-Krum Untargeted-Med
 Frac Data FedAvg Med FLARE FedCC FedAvg Med FLARE FedCC
@@
 The proportion of malicious clients is fixed at 0.3, such that the malicious clients are 3, 9, and 15, respectively.
-Table IV summarizes the accuracy.
-We can observe that the accuracy tends to increase as the participation fraction is greater under Fang-Krum attacks.
+Table 6 summarizes the accuracy.
+We can observe that the accuracy tends to increase as the participation fraction is greater under untargeted-Krum attacks.
 It implies that the more client participation in a non-IID setting, the more accurate the global model is if the attacks are mitigated.
-Under Fang-Med attacks, the accuracy is the highest with FedCC and second highest with Coomed.
+Under untargeted-Med attacks, the accuracy is the highest with FedCC and second highest with Coomed.
 Throughout the experiment, FLARE prevented the global model from convergence, and the test accuracy fluctuated.
 The fluctuation is presumably due to the weight scaling while randomly selecting clients.
-Distinct from previous experiments, the server chooses a fraction of the client every round, meaning not-yet-trained models could be selected.
+Different from previous experiments, the server chooses a fraction of the client every round, meaning not-yet-trained models can be selected.
 In this environment, weight scaling leveraged by FLARE constantly improperly scales the local weights, causing divergence and fluctuation.
@@
 This is due to the rising number of aggregated local models; the more models are aggregated, the more general global model is generated.
-VII-C Impact of the Number of Local Epochs
-TABLE V: Accuracy under Backdoor Attack with doubled numbers of local epochs
-Targeted (Diff Local Epoch)
-Dataset FedAvg Coomed FLARE FedCC
-fMNIST 69.40 72.11 72.32 82.31
-CIFAR10 53.85 49.18 21.61 54.94
-Figure 6: Backdoor confidence when the numbers of local epochs are doubled.
-To examine the impact of the number of local epochs, we doubled up the number of local epochs such that the local models are trained for six epochs in benign clients and twelve epochs in malicious clients.
-The main task accuracy and confidence are summarized and illustrated in Table V and Figure 6.
-As expected, increasing the number of local epochs would not affect the accuracy or confidence but bring the convergence forward while being better at maintaining test accuracy.
-It is true that Coomed and FLARE significantly reduce the backdoor confidence, as well as FedCC; however, the test accuracy has even increased with FedCC due to its precise discrimination against malicious clients.
-VIII Conclusion and Future Work
+7 Conclusion and Future Work
 FL has emerged in response to the rising concerns about privacy breaches while employing AI techniques.
-FL, however, is vulnerable to model poisoning attacks due to a server’s blindness to local datasets, making it challenging to assume data distribution and model parameter integrity.
-In line with this, existing robust aggregation algorithms and defense approaches are orthogonal to various attacks and lack consideration of non-IIDness.
-Throughout exhaustive experiments, FedCC mitigates both untargeted and targeted (or backdoor) attacks while demonstrating its effectiveness in non-IID data environments.
-We leave theoretical guarantees of FedCC and experiments when clients have distinct neural network architectures or tasks as future work.
+FL, however, is vulnerable to model poisoning attacks due to a server’s blindness to local datasets, making it difficult to assume data distribution and model parameter integrity.
+In line with this, many robust aggregation algorithms and defenses have been proposed; still, existing approaches are orthogonal to various attacks and lack consideration of non-IIDness.
+Throughout the exhaustive experiments, FedCC mitigates both untargeted and targeted (or backdoor) attacks while demonstrating its effectiveness in non-IID data environments.
+We leave theoretical guarantees of FedCC and experiments when clients have different neural network architectures or tasks as future work.
 Appendix
-VIII-A Dendrograms to Measure Cluster Distance
-Figure 7: Dendrogram in a non-IID setting
-Figure 7 illustrates the distance between two clusters of clients using dendrograms with a single linkage metric and correlation distance method.
-The first two and last columns correspond to untargeted (Fang-Krum, Fang-Med) and targeted attacks, respectively.
-The total number of clients is set to 10; the participation fraction is 1.0.
-The number of malicious clients is 3 for untargeted and 1 for a targeted attack, with fedAvg.
-Observably, the height difference between the root of clusters is maximum in the second last layer (PLR).
-VIII-B Model Architectures
-The tables below summarize model architectures for each benchmark dataset (fMNIST, CIFAR10, CIFAR100).
-Considering that the end devices are generally incapable of handling heavy computation due to resource or communication constraints, we used lightweight models for the experiments.
-TABLE VI: CNN Architecture for fMNIST dataset
-Layer In Out Ker / Str / Pad Activation
-conv2d_1 1 64 $5\times 5$ / 1 / 0 ReLU
-conv2d_2 64 64 $5\times 5$ / 1 / 0 ReLU
-dropout - - 0.25 -
-flatten - 25600 - -
-fc_1 25600 128 - -
-dropout - - 0.5 -
-fc_2 128 10 - -
-TABLE VII: CNN Architecture for CIFAR10 dataset
-Layer In Out Ker / Str / Pad Activation
-conv2d_1 3 64 $3\times 3$ / 1 / 0 ReLU
-maxpool2d $2\times 2$ / - / 0 -
-conv2d_2 64 64 $3\times 3$ / 1 / 0 ReLU
-maxpool2d $2\times 2$ / - / 0 -
-conv2d_3 64 64 $3\times 3$ / 1 / 0 ReLU
-maxpool2d $2\times 2$ / - / 0 -
-flatten - 256 - -
-dropout - - 0.5 -
-fc_1 256 128 - -
-fc_2 128 10 - -
-TABLE VIII: LeNet Architecture for CIFAR100 dataset
-Layer In Out Ker / Str / Pad Activation
-conv2d_1 3 6 $5\times 5$ / 1 / 0 ReLU
-maxpool2d $2\times 2$ / 2 / 0 -
-conv2d_2 6 16 $5\times 5$ / 1 / 2 ReLU
-maxpool2d $2\times 2$ / 2 / 0 -
-flatten - $4$
-fc_1 44944 120 - ReLU
-fc_2 120 84 - ReLU
-fc_3 84 100 - -
-Acknowledgment
+Appendix 0.A Dataset and Model Architecture
+We use three benchmark vision datasets: Fashion-MNIST (fMNIST) [32], CIFAR10, and CIFAR100 [14].
+fMNIST consists of 60,000 training and 10,000 testing samples of 28$\times$28 grayscale images and associating labels of 10 classes.
+The labels include T-shirt/top, Trouser, Pullover, Dress, Coat, Sandal, Shirt, Sneaker, Bag, and Ankle boot.
+CIFAR10 consists of 50,000 training and 10,000 testing examples of 32$\times$32 color images and associating labels of 10 classes.
+The classes include airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck.
+CIFAR100 is the same as CIFAR10 except for the categorization.
+It is labeled as 100 classes, each containing 600 images.
+Thus, there are 500 training and 100 testing images per class.
+Each class (i.e., ‘fine’ label) belongs to one of the 20 superclasses.
+Appendix 0.B FedCC: Algorithm
+Algorithm 1 summarizes FedCC’s workflow.
+Algorithm 1 FedCC
+Input: global_w, n local_w ($w_{1}$, …, $w_{n}$)
+      Output: agg_w, larger_cluster_members
+$\triangleright$ Get PLRs of local models
+1:for $i<n$ do
+2:     local_plrs[$i$] $\leftarrow$ local_w[$i$][second_last_layer]
+3:end $\triangleright$ Extract global PLR
+4:glob_plr $\leftarrow$ global_w[second_last_layer] $\triangleright$ Apply kernel CKA to the plrs
+5:cka[$i$] $\leftarrow$ kernel_CKA(glob_plr, local_plrs[$i$]) $\triangleright$ Normalize similarities globally
+6:across_cka $\leftarrow$ normalize(cka) $\triangleright$ Apply Kmeans clustering algorithm
+7:kmeans $\leftarrow$ kmeans(n_clusters=2, cka)
+8:labels $\leftarrow$ kmeans.labels_
+9:count $\leftarrow$ counter(labels)
+10:larger_cluster $\leftarrow$ 1
+11:if count[00] $>$ count[$1$then
+12:     larger_cluster $\leftarrow$ 0
+13:end $\triangleright$ Identify larger cluster members
+14:larger_cluster_members = where(labels == larger_cluster) $\triangleright$ Normalize similarities within larger cluster
+15:within_cka $\leftarrow$ normalize(cka[larger_cluster_members])
+16:Initialize agg_weights as zero tensor $\triangleright$ Differently weigh weights per layer
+17:for layer in model_layers do
+18:     if layer is up to second-to-last layer then
+19:         agg_weights[layer] $\leftarrow$ weighted average using across_cka
+20:     else if layer is second-to-last layer then
+21:         agg_weights[layer] $\leftarrow$ weighted average using within_cka
+22:     else
+23:         agg_weights[layer] $\leftarrow$ weighted average using across_cka
+24:     end if
+25:end for
+26:return agg_weights, larger_cluster_members

### v4

@@
 However, it introduces new attack surfaces, which are especially prone when data is non-Independently and Identically Distributed.
-Existing approaches fail to effectively mitigate the malicious influence in this setting; previous approaches often tackle non-IID data and poisoning attacks separately.
+Previous approaches often tackle non-IID data and poisoning attacks separately.
 To address both challenges simultaneously, we present FedCC, a simple yet effective novel defense algorithm against model poisoning attacks.
@@
 We believe that this new perspective on aggregation makes it a valuable contribution to the field of FL model security and privacy.
-The code will be made available upon acceptance.
+Code is available at https://github.com/HyejunJeong/FedCC.
 Keywords:
@@
 1 Introduction
-Federated Learning (FL) [22] is a distributed model training framework designed to preserve privacy.
-FL allows data to remain on client devices but exchanges their local model parameters only for model training.
-It thus reduces (data) privacy risks associated with centralized learning, not requiring central data collection.
-Thus, FL is especially suitable for data-sensitive environments since raw data remains on clients, reducing the risk of data leakage on the server or in between communication.
+Federated Learning (FL) [23] is a distributed model training framework designed to preserve privacy by restricting data to remain on client devices.
+Only model parameters are exchanged, minimizing data privacy risks typically associated with centralized learning.
+This makes FL particularly useful in data-sensitive environments, where raw data should never leave clients, reducing the risk of data leakage.
 However, its distributed nature makes FL susceptible to model poisoning attacks [13].
@@
 Backdoor attacks are stealthier targeted attacks that maintain overall performance while misclassifying specific inputs [1].
-Some preliminary defenses propose robust aggregation methods, but they may compromise privacy by sharing raw data or exposing data distribution to untrusted parties [6, 30].
-While some defenses exist, most focus on IID data settings, which are rare in real-world scenarios.
-Non-IID data, with variations in class distribution and features across clients [12], can gradually degrade model performance and complicate the detection of malicious clients [39, 35], as the impact of attacks increases with the degree of non-IID [19, 26].
-In this study, we introduce FedCC, a defense mechanism that mitigates untargeted model poisoning and targeted backdoor attacks in both IID and non-IID settings.
-FedCC leverages the Centered Kernel Alignment (CKA) of Penultimate Layer Representations (PLRs) to distinguish malicious clients from benign ones.
-PLRs are known to be highly specific to local data distributions [36], making clients distinguishable, especially when data is non-IID.
-Exploiting CKA’s well-known capability of measuring similarity in high-dimensional space, such as neural network representations, we accurately measure the similarity between the PLR of different clients, further allowing more accurate detection of malicious clients.
-Empirical evidence supporting our rationale is shown in Figure 2 and Table 1.
-The use of PLRs shows the highest separability between benign and malicious clients in all settings, as seen by the largest distance between clusters in the PLR.
-The use of CKA, specifically kernel CKA, is justified by its highest performance in both targeted and untargeted IID and non-IID settings.
-Furthermore, our method maintains data privacy by using similarity measures between local and global models, not having shared data.
-We named our proposed method FedCC, Federated learning using Clustered CKA.
-Empirical experiments show that FedCC outperforms existing defenses that rely on first-order statistics, which usually suffer from a high false negative detection rate.
-It significantly improves global accuracy against untargeted attacks (66.27% compared to similar approaches) and reduces attack confidence in targeted backdoor attacks to almost zero while preserving main task accuracy.
-These results highlight FedCC’s effectiveness and robustness in mitigating attacks in FL settings, particularly in non-IID data settings.
+While some defenses rely on robust aggregation methods, many fail to maintain privacy by sharing raw data or exposing data distributions to untrusted parties [7, 31].
+Furthermore, most existing defenses assume IID (Independent and Identically Distributed) data, which is rare in real-world scenarios.
+Non-IID data, where clients’ data distributions and features vary, complicates attack detection and performance maintenance [13, 39, 36].
+As the degree of non-IID data increases, the impact of attacks also grows [20, 27].
+In this study, we introduce FedCC, a defense mechanism against untargeted model poisoning and targeted backdoor attacks in both IID and non-IID settings.
+FedCC leverages Centered Kernel Alignment (CKA) of Penultimate Layer Representations (PLRs) to distinguish malicious clients from benign ones.
+PLRs are highly sensitive to local data distributions, making clients distinguishable, especially in non-IID environments.
+By exploiting CKA’s ability to measure similarity in high-dimensional spaces, we can accurately identify malicious clients.
+Empirical evidence, shown in Figure 3 and Table 1, demonstrates that PLRs provide the highest separability between benign and malicious clients, with CKA delivering superior performance in both IID and non-IID settings.
+Importantly, FedCC ensures data privacy by relying on similarity measures between models instead of shared data.
+The proposed method outperforms existing defenses based on first-order statistics, which typically suffer from high false negative rates.
+Our experiments show that FedCC significantly improves global accuracy against untargeted attacks (66.27% improvement) and reduces attack confidence in targeted backdoor attacks to nearly zero while preserving main task accuracy.
 Our contributions are summarized as follows:
-  • We propose FedCC, a novel and scalable defense method against model poisoning attacks using the CKA of PLRs of local and global models and performing layer-wise weighted aggregation of their model parameters.
-  • We justify that PLR is a highly distinguishable feature to detect malicious clients by measuring the discrepancy between two clusters of clients.
-  • We demonstrate that CKA provides the most sensitive and accurate similarity measure between differently trained models when the server has no knowledge about clients or centralized data, even in non-IID settings.
-  • We empirically demonstrate the effectiveness of FedCC through exhaustive experiments.
-Our results show that comparing the CKA of PLRs filters malicious clients more reliably than the first-order statistics.
-FedCC outperforms existing defenses even in non-IID settings.
+  • We propose FedCC, a novel and scalable defense method that uses CKA of PLRs and performs layer-wise weighted aggregation of model parameters to defend against model poisoning attacks.
+  • We demonstrate that PLRs provide a highly distinguishable feature for detecting malicious clients by measuring the discrepancy between clusters of clients.
+  • We justify the use of CKA as an accurate and sensitive similarity measure for comparing models, even in non-IID settings, when the server has no access to client data.
+  • We empirically validate the effectiveness of FedCC through extensive experiments, showing that it outperforms existing defenses, especially in non-IID scenarios.
 2 Backgrounds and Related Works
@@
 2.3 Penultimate Layer Representation in FL
-The penultimate layer is a neural network model’s second-to-last layer (i.e., before the softmax layer).
+Figure 1: LeNet (CNN) architecture and Penultimate Layer
+The penultimate layer is a neural network model’s second-to-last layer (i.e., before the softmax layer), as shown in Figure 1.
 Wang et al.
-[30] discovered that PLRs are effective in distinguishing malicious models from benign ones; within benign clients, benign PLRs follow the same distribution, whereas malicious PLRs do not, across datasets and neural network architectures.
-Specifically, they demonstrated that distances among benign PLRs are smaller than those of benign and malicious PLRs.
+[31] discovered that PLRs are effective in distinguishing malicious models from benign ones; within benign clients, PLRs follow the same distribution, whereas malicious PLRs do not—across datasets and neural network architectures.
+Specifically, they demonstrated that distances among benign PLRs are smaller than those between benign and malicious PLRs.
 Their proposed method, FLARE, assigned a trust score to each client based on pairwise PLR discrepancies defined by Maximum Mean Discrepancy among all model updates, allocating lower values to those farther from the benign distribution.
 The model updates are then scaled and averaged, weighted by each client’s trust score.
-Meanwhile, FLARE continuously redirects the global model using server-owned raw data, increasing the risk of data leakage and jeopardizing data privacy.
+Meanwhile, FLARE continuously redirected the global model using server-owned raw data, increasing the risk of data leakage and jeopardizing data privacy.
 2.4 Centered Kernel Alignment
-Centered Kernel Alignment (CKA) [13] is a highly accurate similarity metric that measures how similar two differently initialized or trained neural networks are, producing a value between 0 (not similar) and 1 (identical).
-CKA can be calculated by: $CKA(X,Y)=\frac{HSIC(X,Y)}{\sqrt{HSIC(X,X)HSIC(Y,Y)}}=\frac{{\lVert K(X^{T})K(Y)\rVert}^{2}}{{\lVert K(X^{T})K(X)\rVert}{\lVert K(Y^{T})K(Y)\rVert}}$ where $HSIC$ is Hilbert-Schmidt Independence Criterion [9] (i.e., a not normalized version of CKA) and $K$ is the RBF kernel.
-It constructs similarity kernel matrices in the weight space and compares these matrices to characterize the representation space, allowing comparisons between layers with varying widths and different initialization schemes.
-CKA includes several vital properties for similarity measures of neural networks.
+CKA [14] is a highly accurate similarity metric used to measure how similar two differently initialized or trained neural networks are, producing a value between 0 (no similarity) and 1 (identical).
+It is computed as:
+$CKA(X,Y)=\frac{HSIC(X,Y)}{\sqrt{HSIC(X,X)HSIC(Y,Y)}}=\frac{{\lVert K(X^{T})K(Y)\rVert}^{2}}{{\lVert K(X^{T})K(X)\rVert}{\lVert K(Y^{T})K(Y)\rVert}}$
+where $HSIC$ refers to the Hilbert-Schmidt Independence Criterion [10], a non-normalized variant of CKA, and $K$ denotes the RBF kernel.
+CKA constructs similarity kernel matrices in the weight space and compares them to characterize the representation space, enabling comparison across layers with different widths and initialization schemes.
+CKA satisfies several key properties desirable in neural network similarity metrics.
 It is non-invariant to invertible linear transformations, meaning similarity scores change if such transformations are applied.
 This property is vital; the gradient descent algorithm is not invariant to these transformations because similarity metrics invariant to linear transformations are inaccurate on models trained with gradient descent.
-It is also invariant to orthogonal transformations and isotropic scaling, which are desirable for neural networks trained with gradient descent owing to their stochastic nature.
-These properties make CKA a precise and reliable metric for measuring similarities between neural networks, particularly when comparing potential models from malicious clients.
-CKA has been utilized to address data heterogeneity issues, taking into account that the similarity of non-IID data is notably lower than that of IID data [28], particularly in specific layers [27].
-However, CKA has yet to be leveraged as a defense mechanism against attacks involving models from malicious clients where the data are distributed across the clients in a non-IID manner.
+It is also invariant to orthogonal transformations and isotropic scaling, both of which are desirable for neural networks trained with gradient descent owing to their stochastic nature.
+These properties make CKA a precise and reliable metric for measuring similarities between neural networks, particularly when comparing models potentially generated by malicious clients.
+CKA has been utilized to address data heterogeneity issues, taking into account that the similarity of non-IID data is notably lower than that of IID data [29], especially in specific network layers [28].
+However, it has not yet been explored as a defense mechanism against malicious clients, where model similarity could reveal suspicious deviations under non-IID distributions.
 3 Threat Model
@@
   • An attacker has full control over at most $k<n/2$ clients.
-  • An attacker has knowledge about compromised clients’ data, such as current local model, previous global model, and hyperparameters.
+  • An attacker has knowledge about compromised clients’ data, such as the current local model, previous global model, and hyperparameters.
   • An attacker has no knowledge or control over the server and honest clients.
@@
 Untargeted-Med [8] manipulates the model parameters based on the maximum and minimum so that chosen coordinate-wise median values direct toward an inverse direction.
-The attack starts with defining $w_{max,j}=\text{max}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$, which indicates the maximum, and $w_{min,j}=\text{min}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$ which indicates the minimum of the $j$th local model parameters on the benign clients, respectively.
+The attack starts with defining the maximum and minimum of the $j$th local model parameters on the benign clients, $w_{max,j}=\text{max}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$ and $w_{min,j}=\text{min}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$, respectively.
 Also, to avoid sampled $m$ numbers being outliers, if $s_{j}=-1$, $m$ numbers in [$w_{max,j},b\cdot w_{max,j}$] (when $w_{max,j}>0$) or [$w_{max,j},b/w_{max,j}$] (when $w_{max,j}\leq 0$) are randomly sampled.
@@
 For DBA [34], we implant split backdoor triggers into the input so that the global model can be attacked in a distributed manner.
-For more detail, refer to [33].
-In CIFAR10, for example, we manipulate the model parameters to misclassify an image of ‘airplanes’ as ‘bird’ whereas normally classifying the image of other classes.
+For more details, refer to [34].
+In CIFAR10, for example, we manipulate the model parameters to misclassify an image of ‘airplanes’ as ‘birds’, whereas normally it would classify the image as other classes.
+Figure 2: An Overview of FedCC.
+$k$, $n$, $a$, and $b$ indicate the kth layer, the number of participating clients, normalized similarity across all layers (i.e., across_cka), and normalized similarity within the cluster (i.e., within_cka), respectively.
 4 FedCC: Robust Aggregation against Poisoning Attacks
@@
 Figure 2 depicts an overview of FedCC.
-(1) A server initializes and advertises the global model to $n$ participating clients, selected out of total $k$ clients with a probability of fraction $C$.
-(2) The clients individually train their local models using their respective datasets and send the trained weights to the server for aggregation.
-Our proposed method intervenes at this moment.
+(1) The server initializes and broadcasts the global model to $n$ clients, selected from a total of $k$ candidates with selection fraction $C$.
+(2) Each client trains a local model on its private dataset and sends the updated weights to the server—this is where our proposed method intervenes.
 (3) The server extracts PLRs from both the global model and each local model, and (4) compares RBF Kernel CKA values between them.
 (5) The CKA values are then clustered into two, potentially representing benign and malicious clients.
-(6) The CKA similarity values are then normalized across all clients for every layer as $a$, but the second last layer’s CKA similarity values are normalized within the larger cluster as $b$, backed by the assumption that the number of malicious clients does not exceed $n/2-1$.
+(6) CKA similarity values are normalized across all clients for each layer as $a$.
+For the second-to-last layer (PLR), normalization is performed within the larger cluster as $b$, under the assumption that malicious clients are fewer than $n/2-1$.
 (7) Finally, the server aggregates each layer of local models weighted by $a$, but the PLR weighted by $b$, to update a global model for the current epoch and proceeds to the next epoch by distributing the updated global model to the newly selected $n$ clients.
-Figure 1: An Overview of FedCC.
-$k$, $n$, $a$, and $b$ indicate the kth layer, the number of participating clients, normalized similarity across all layers (i.e., across_cka), and normalized similarity within the cluster (i.e., within_cka), respectively.
 4.2 Detailed Design
-A major challenge when filtering malicious clients in the presence of non-IID data is the server’s lack of knowledge regarding the data distribution.
+A key challenge in filtering malicious clients under non-IID conditions is the server’s ignorance of the underlying data distribution.
 Even if the data is benign, local learning models can exhibit significant angular or magnitude differences if they are non-IID.
@@
 In federated learning, where the server cannot access local training data, it receives trained model parameters (weights) from selected clients.
-Therefore, the collected weights are the only available information for determining whether a client is malicious.
-PLR takes the end result of the convolution and pooling process that will reach a classification decision.
-Thus, we assumed that computed weights of the second last layer might contain the essential information that strongly influences classification results.
+Thus, the received weights serve as the sole basis for detecting malicious behavior.
+The PLR captures the final representation produced by convolution and pooling before reaching a classification decision.
+We hypothesize that the PLR contains the most task-relevant, discriminative features.
 Figure 3: Cluster Distance of Each Layer
-To justify the utilization of PLRs, we measure the distance between two clusters of clients (benign and malicious) using dendrograms with a single linkage metric and correlation distance method.
+To test this, we measure the distance between two clusters of clients (benign and malicious) using dendrograms with a single linkage metric and correlation distance method.
 Figure 3 presents cluster distances of each layer under different attack scenarios, such as untargeted-mKrum attacks in IID or non-IID settings.
@@
 Therefore, we utilize the weights of the penultimate layer to compare their similarity to the global model.
-4.2.2 Kernel CKA[13]
-is a similarity metric to assess the similarity or dissimilarity between two neural networks.
-It compares representations at different layers or models trained in different ways.
-The sophisticated similarity comparison ability of Kernel CKA makes it a suitable choice for our purpose.
-To justify the effectiveness of Kernel CKA, we compared its performance with other similarity metrics such as linear CKA, MMD (not normalized CKA), cosine similarity (angle), and Euclidean similarity (distance).
-We evaluated the test accuracy of FedCC using the CIFAR10 dataset under two untargeted attacks and one targeted attack, separately in IID and non-IID settings.
-The results are summarized in Table 1.
-In the ’targeted’ column, the values represent the combination of test accuracy and backdoor confidence, with higher test accuracy and lower backdoor confidence indicating better performance.
-We observed that Kernel CKA consistently yields the highest performance across experiments; therefore, we decided to utilize Kernel CKA to compare the similarities of PLRs.
 Table 1: Comparison of Performance with Various Similarity Metrics
@@
 Euclidean 69.06 27.82 68.54 41.57 69.17/0.0221 52.20/0.0015
+4.2.2 Kernel CKA[14]
+is a similarity metric to assess the similarity or dissimilarity between two neural networks.
+It enables the comparison of representations across layers or among models trained under varying conditions.
+The sophisticated similarity comparison ability of Kernel CKA makes it suitable for our purpose.
+To justify the effectiveness of Kernel CKA, we compared its performance with other similarity metrics such as linear CKA, MMD (not normalized CKA), cosine similarity (angle), and Euclidean similarity (distance).
+We evaluated the test accuracy of FedCC using the CIFAR10 dataset under two untargeted attacks and one targeted attack, separately in IID and non-IID settings.
+The results are summarized in Table 1.
+In the ‘targeted’ column, the values represent the combination of test accuracy and backdoor confidence, with higher test accuracy and lower backdoor confidence indicating better performance.
+We observed that Kernel CKA consistently yields the highest performance across experiments; therefore, we adopted it to measure PLR similarity.
 4.2.3 FedCC
-is an aggregation method that combines local clients’ model parameters, based on the Kernel CKA similarities between the global and local models’ PLRs.
-An overall algorithm is described in Appendix 0.B.
-A server first extracts PLRs of the global and local models and calculates the kernel CKA between them.
-We specifically focus on comparing PLRs because they exhibit the most distinguishable differences among differently trained models, as demonstrated in previous sections.
-Additionally, we utilize RBF Kernel CKA, as it effectively captures similarity differences between non-IID and malicious clients.
-The similarity values obtained are then clustered using a simple K-means clustering algorithm.
-Any clustering method that produces only two clusters can be used.
-Considering the assumption that the number of attackers is less than half of the total number of clients ($n/2$), we designate the clients in the larger cluster as ‘candidates.’
-Since the earlier layers in CNNs learn global features and the latter layers embed local features [15, 25, 23], we used two differently normalized CKA similarity values: one normalized within the candidates (i.e., within_cka, denoted as $b$ in Figure 1) and one normalized across all the clients (i.e., across_cka, denoted as $a$ in Figure 1).
-For the PLR (second-last layer), we apply within_cka to the clients belonging to the larger cluster, whereas for all other layers, we apply across_cka to weighted average model parameters.
-By doing so, we can ensure that the end magnitude of the aggregated parameters stays similar to before the manipulations, leveraging the fact that the sum of the normalized values is one.
-Moreover, since our method evaluates clients in every global round and does not exclude any local models, it is also effective for adaptive attacks like DBA.
-5 Experiments
-5.1 Dataset and Model Architecture
-We use three benchmark vision datasets: Fashion-MNIST (fMNIST) [32], CIFAR10 and CIFAR100 [14].
-Details about datasets can be found in Appendix 0.A.
-Considering that the end devices are normally incapable of handling heavy computation due to resource or communication constraints, we used lightweight CNN architectures for the experiments.
-For fMNIST, we used a 4-layer-CNN with two convolutions and two fully connected (FC) layers, incorporating dropout with probabilities of 0.25 and 0.5 before and after the FC layer, respectively.
-For CIFAR10, we utilized a 5-layer-CNN with three convolutions, each followed by a max-pool layer and two FC layers, applying dropout with a probability of 0.5 before the first FC layer.
-For CIFAR100, we adopted LeNet [16].
-5.2 Non-IID Simulation
-Non-IID is a more feasible assumption considering the diverse and massive nature of data in practice.
-To standardize non-IID settings, we use Dirichlet distribution, which is widely used and well-known for its capability of representing real-world data [18].
-A concentration parameter $\alpha$ controls the degree of non-IID; the smaller $\alpha$ is, the more likely the clients hold examples from only one randomly chosen class.
-We set the $\alpha$ to 0.2 to represent non-IID, based on [11, 38].
-5.3 Experimental Setup and Baselines
-We implement two untargeted attacks from [7] and two targeted attacks from [3] and [33], along with defense baselines including Krum, multi-Krum [4], Coomed [34], Bulyan [10], FLTrust [6], FLARE [30], and FedCC.
-We use ten clients with a 1.0 client participation fraction unless stated otherwise.
-As in [30], we have three adversaries for each untargeted attack and one for the targeted attack.
-In the DBA, we randomly select four malicious clients and implant five distributed triggers per batch.
-To compensate for delayed attack effects, we use FedAvg until epoch 30, after which each defense method is applied.
-Note that, for a fair comparison, FLTrust is computed using its previous global model without auxiliary data.
-Benign clients undergo three local epochs, while compromised clients undergo six, with training done using the Adam optimizer (learning rate = 0.001).
-5.4 Evaluation Metrics
-We evaluated our defense mechanism using two metrics: backdoor confidence (confidence) and global model test accuracy (accuracy).
-The confidence metric measures misclassification likelihood, ranging from 0 to 1 (lower values are better).
-Accuracy reflects the global model’s overall test performance under defense methods during attacks, ranging from 0 to 100 (higher values are better).
-Both metrics were reported under backdoor attacks, with a focus on accuracy for untargeted attacks, as the attacker’s goal is to reduce overall accuracy.
-6 Results and Discussion
-In this section, we report the performance of our FedCC against two untargeted and two targeted attacks in both IID and non-IID environments.
-We additionally demonstrate the defense effect against a distributed adaptive attack, DBA [33].
-Unlike centralized backdoor attacks, the DBA splits a trigger pattern across multiple adversaries, evading the detection by anomaly detection.
-We show the result of CIFAR10, with various defense methods applied.
-6.1 Non-IID Data Environment
-Non-IID is a more feasible assumption considering the diverse and massive nature of data in practice.
-To standardize non-IID settings, we employ Dirichlet distribution that 0 indicates the most heterogeneity, and 100 mimics homogeneity [20].
-A concentration parameter $\alpha$ controls the degree of non-IID; the smaller $\alpha$ is, the more likely the clients hold examples from only one randomly chosen class.
-Since $\alpha$ being 0.2 represents a highly non-IID scenario based on [11, 38], we set it as such, following [38].
-Table 2: Test Accuracy under untargeted attacks in Non-IID setting.
-Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
-Fang -Krum non-IID fM 64.63 16.51 45.91 57.12 13.39 49.37 49.54 66.27
-C10 33.69 15.38 20.09 35.70 19.23 36.63 17.03 37.40
-C100 2.27 1.00 5.34 4.85 0.98 11.62 7.46 14.51
-Fang -Med non-IID fM 16.32 49.33 66.30 68.90 64.12 12.38 52.25 72.76
-C10 10.02 25.06 40.52 40.23 32.47 12.39 14.59 42.37
-C100 1.00 6.24 13.97 10.27 6.91 1.00 1.00 16.12
-6.1.1 Untargeted Model Poisoning Attacks in non-IID setting
-Table 2 provides the global model’s accuracy across three datasets under two untargeted (Fang-Krum and Fang-Med) in a non-IID environment.
-A notable finding is that robust aggregation algorithms often yield lower accuracy than simple averaging (fedAvg).
-This discrepancy arises from the imperfect identification of malicious clients, leading to the erroneous aggregation of their weights along with those of benign clients.
-Under Fang-Krum, Krum, Bulyan, and multi-Krum experience significant accuracy drops (16.51%, 13.39%, and 45.91%, respectively, compared to FedAvg’s 64.63%), reflecting the attack’s design to degrade Krum-based defenses.
-Coomed also struggles, as median weights fail to represent benign clients effectively in non-IID settings.
-FLTrust mitigates Fang-Krum but performs better on CIFAR10 than fMNIST, likely due to fMNIST’s simpler gradients and lower variance, which challenge its ability to distinguish malicious updates.
-FLARE’s lower accuracy stems from its approach of scaling the entire model weights uniformly, which ignores knowledge contributed by individual clients.
-In contrast, FedCC, which selectively averages weights layer-wise, achieves superior performance by boosting or minimizing weights based on their alignment with benign updates.
-FedCC demonstrates the highest accuracy (66.27%, 37.40%, and 14.51% for fMNIST, CIFAR10, and CIFAR100, respectively) and shows significant improvement in CIFAR100 experiments.
-Turning to the Fang-Med attack, FedAvg experiences a significant drop in accuracy due to the attack’s creation of outliers by deviating from median values.
-Consequently, Coomed and Bulyan perform relatively well since their coordinate-wise median and trimmed mean methods disregard outliers effectively.
-FLARE continues to perform poorly for a similar reason mentioned earlier.
-Between the Krum-based methods, multi-Krum and Coomed are more robust than Krum, as they aggregate multiple local models, providing better resilience to outliers.
-Similarly, Bulyan’s averaging of multiple clients enhances robustness but falls short of Coomed, particularly in non-IID environments.
-Finally, FedCC achieves the highest accuracy (72.76%, 42.37%, and 16.12% for each dataset) without sharing raw data, showcasing its exceptional performance.
-This success is attributed to two factors: (1) highly distinguishable information within PLRs and (2) higher CKA similarity between benign clients than between benign and malicious clients.
-6.1.2 Targeted Backdoor Attacks
-Table 3: Test Accuracy under targeted attacks in Non-IID setting.
-Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
-Target non-IID fM 75.65 45.27 65.97 71.70 57.96 61.82 64.31 75.66
-C10 36.16 14.98 30.72 48.97 40.11 44.06 10.18 51.56
-C100 4.46 6.18 6.90 12.04 11.16 12.95 1.14 15.26
-DBA C10 38.56 24.94 7.09 44.45 34.19 51.49 38.73 52.28
-Figure 3: Confidence of Backdoor Task for targeted attacks in Non-IID settings.
-Table 3 summarizes test accuracy under targeted attacks in a non-IID setting.
-To further illustrate the impact, Figure 3 provides a visual representation of backdoor confidence.
-It is evident that Krum, which selects a single client’s weights as the global model, shows reduced robustness compared to methods that aggregate weights from multiple clients.
-In contrast, Coomed, multi-Krum, and Bulyan exhibit higher test accuracy.
-FLARE reduces main task accuracy for reasons similar to those observed in previous experiments.
-FedCC achieves the highest test accuracy.
-Since targeted attacks aim to misclassify a specific target class while maintaining the main task accuracy, the best performance of FedCC is attributed to leveraging knowledge from earlier stages of training.
-This result highlights the superiority of using sophisticated weighted knowledge and filtering malicious clients via CKA similarity, which outperforms first-order statistical methods like mean or median.
-A similar trend is observed with the DBA attack, except for multi-Krum; the lowest accuracy occurs because all four malicious clients, each with a distributed trigger, are treated as outliers and excluded from aggregation.
-Notably, Figure 3 demonstrates that FedCC significantly reduces backdoor confidence.
-Note that since the backdoor confidence of DBA fluctuates a lot due to its distributed nature, we omit it for brevity.
-Unlike other methods, such as FLTrust or multi-Krum, which exhibit fluctuating confidence values, FedCC maintains consistently low confidence, emphasizing its resilience and independence from client-specific data distributions.
-This advantage stems from the effective utilization of CKA, enabling similarity calculations even when models are trained on different datasets with varying distributions.
-While other baselines, like Bulyan or FLTrust, also reduce confidence, we argue that FedCC not only preserves but potentially increases accuracy by precisely distinguishing malicious clients.
-6.2 IID Data Environment
-In an IID setting, we evenly divide the training dataset among all clients so that each client’s data distribution is identical and of the same size.
-Table 4: Test Accuracy under untargeted attacks in IID setting.
-Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
-Fang -Krum IID fM 75.55 31.66 87.78 87.62 50.30 89.53 79.16 89.57
-C10 49.67 40.86 63.42 57.40 12.67 68.25 25.77 69.84
-C100 13.72 1.04 7.64 6.17 1.59 17.14 7.49 18.47
-Fang -Med IID fM 20.86 85.33 89.53 86.70 87.45 21.36 71.08 89.66
-C10 9.51 54.28 69.68 59.20 57.69 9.92 49.82 70.52
-C100 0.87 12.27 16.52 14.43 12.56 1.16 5.83 17.83
-6.2.1 Untargeted Model Poisoning Attacks
-Table 4 shows the test accuracy of the global model trained on three datasets under untargeted attacks in IID settings.
-While trends in these experiments are similar to those in the non-IID setting, the accuracy values are slightly higher due to the more consistent data distribution across and updates from clients in the IID case.
-A notable deviation from the non-IID experiments is when Coomed is applied: the accuracy increases from 75.55% (FedAvg) to 87.62% (Coomed).
-This indicates that the coordinate-wise median values closely align with the global model’s weights with minimal deviation and represent benign models well, as the adversary’s impact is limited due to the majority of benign clients.
-FedCC achieves the highest accuracy across all methods (89.57%, 69.84%, and 18.47% for each dataset) due to its effective CKA similarity measurement and sophisticated layer-wise aggregation strategy.
-Under the Fang-Med attack, the accuracy of FedAvg experiences a significant drop (89.89% to 20.86%) due to the peculiarity of outliers when the clients have IID data.
-Similar to the result of Fang-Krum, Krum, Coomed, and Bulyan demonstrate similar accuracy values, indicating that coordinate-wise median values indeed represent benign clients’ parameters.
-FLTrust, however, is less effective in mitigating the Fang-Med attack than Fang-Krum, even in the IID setting; it then means that the performance difference is not due to data distribution but rather its inherent vulnerability.
-In Fang-Krum, malicious updates are deliberately far from the global model, allowing FLTrust to identify discrepancies easily.
-In contrast, the malicious updates in Fang-Med are subtler and only slightly deviate from the global model or median, making it harder for FLTrust to distinguish them from benign updates.
-Meanwhile, FedCC consistently achieves the highest accuracy (89.66%, 70.52%, and 17.83% for each dataset), demonstrating its capability to mitigate attacks involving more nuanced changes to model updates without extreme outliers.
-6.2.2 Targeted Backdoor Attacks
-Table 5: Test Accuracy under targeted attacks in IID setting.
-Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
-Target IID fM 88.27 86.63 87.03 89.41 89.45 89.59 75.29 90.01
-C10 64.68 57.69 71.19 69.85 68.76 68.61 11.09 71.64
-C100 13.83 5.72 17.57 13.24 15.11 17.08 1.03 18.61
-DBA C10 10.00 35.58 51.40 10.00 16.16 56.69 10.00 58.04
-Figure 4: Confidence of Backdoor Task for targeted attacks in IID settings.
-Table 5 and Figure 4 present the test accuracy and backdoor confidence under targeted backdoor attacks in IID settings.
-Similar to the untargeted attack scenario, the general test accuracy in IID settings is higher than in non-IID settings.
-Reducing backdoor confidence is as critical as maintaining main task accuracy, as attackers aim to stealthily embed backdoor tasks.
-A key observation is FedCC’s remarkable ability to reduce backdoor confidence to near zero while preserving high main task accuracy.
-In contrast, the fluctuating backdoor confidence observed in Multi-Krum and Flare is largely due to the distributed trigger, which confuses their defense mechanisms and hampers their ability to filter out outliers effectively.
-Other methods, such as Multi-Krum, Flare, and FLTrust, show significant variability in backdoor confidence, further highlighting FedCC’s superiority.
-Notably, FedCC achieved the highest main task accuracy across both targeted backdoor attacks and distributed backdoor attacks (DBA), consistently neutralizing the backdoor task.
-These results underscore FedCC’s ability to mitigate targeted backdoor attacks through precise and efficient integration of prior knowledge to preserve main task performance while zeroing out the attack confidence.
-6.3 Performance in Varying FL Settings
-We additionally measured the effectiveness of FedCC in various non-IID settings, including different numbers of malicious clients, fractions of participation, and numbers of local epochs.
-Since Coomed was as effective as ours, and FLARE also used PLRs, we compared three methods to FedCC, including FedAvg.
-6.3.1 Impact of the Number of Malicious Clients
-Figure 5: Test Accuracy with different numbers of malicious clients
-We assessed the impact of different numbers of malicious clients under untargeted attacks, specifically Untargeted-Krum and Untargeted-Med.
-In this set of experiments, we kept the total number of clients and participation fraction fixed at 10 and 1, respectively.
-To adhere to the assumption that the number of attackers is less than half of the participating clients, we investigated the test accuracy considering scenarios involving up to four malicious clients.
-Figure 5 illustrates the accuracy in the given experimental settings, and it is evident that FedCC outperforms the other defense methods.
-We observe a general trend where the accuracy of defense methods decreases as the number of attackers increases.
-Notably, Coomed experiences the most significant drop in accuracy when transitioning from one malicious client to four malicious clients.
-This can be attributed to the fact that as the participation of malicious clients increases, there is a higher likelihood of the median values being influenced by their malicious contributions.
-In contrast, FedCC does not rely solely on geometric measures such as angles or distances but instead leverages hidden correlations between networks to identify and filter out malicious clients.
-As a result, FedCC demonstrates superior accuracy in filtering out malicious clients, regardless of the number of attackers involved.
-It is important to note that the accuracy degradation observed as the number of malicious clients increases primarily due to aggregating fewer clients rather than the malicious clients themselves being selected.
-These findings underscore the robustness and effectiveness of FedCC in defending against untargeted attacks, as it consistently outperforms other defenses across varying numbers of malicious clients.
-6.3.2 Impact of Fraction
-Table 6: Test Accuracy under untargeted attacks with different fractions of clients being selected.
-Untargeted-Krum Untargeted-Med
-Frac Data FedAvg Med FLARE FedCC FedAvg Med FLARE FedCC
-0.1 fM 55.31 49.83 34.02 64.83 16.57 66.41 52.24 69.52
-C10 10.06 22.55 14.50 20.49 10.00 15.33 10.00 29.81
-0.3 fM 64.22 57.52 10.00 73.55 16.26 58.07 10.00 61.12
-C10 24.24 12.59 10.00 27.81 10.98 22.61 10.00 38.27
-0.5 fM 62.37 58.20 10.00 76.41 18.36 62.49 10.00 69.05
-C10 23.99 17.28 10.06 34.32 9.87 27.83 10.00 37.27
-We examine the impact of the participating fraction of clients where the number of total clients is fixed at 100.
-The fractions are 0.1, 0.3, and 0.5, such that the numbers of clients being aggregated are 10, 30, and 50, respectively.
-The proportion of malicious clients is fixed at 0.3, such that the malicious clients are 3, 9, and 15, respectively.
-Table 6 summarizes the accuracy.
-We can observe that the accuracy tends to increase as the participation fraction is greater under untargeted-Krum attacks.
-It implies that the more client participation in a non-IID setting, the more accurate the global model is if the attacks are mitigated.
-Under untargeted-Med attacks, the accuracy is the highest with FedCC and second highest with Coomed.
-Throughout the experiment, FLARE prevented the global model from convergence, and the test accuracy fluctuated.
-The fluctuation is presumably due to the weight scaling while randomly selecting clients.
-Different from previous experiments, the server chooses a fraction of the client every round, meaning not-yet-trained models can be selected.
-In this environment, weight scaling leveraged by FLARE constantly improperly scales the local weights, causing divergence and fluctuation.
-These experimental results indicate that selectively choosing clients based on CKA similarity is more reliable than taking the median value of each coordinate in a non-IID setting.
-We also observed that as the participation fraction grows, the accuracy grows.
-This is due to the rising number of aggregated local models; the more models are aggregated, the more general global model is generated.
-7 Conclusion and Future Work
-FL has emerged in response to the rising concerns about privacy breaches while employing AI techniques.
-FL, however, is vulnerable to model poisoning attacks due to a server’s blindness to local datasets, making it difficult to assume data distribution and model parameter integrity.
-In line with this, many robust aggregation algorithms and defenses have been proposed; still, existing approaches are orthogonal to various attacks and lack consideration of non-IIDness.
-Throughout the exhaustive experiments, FedCC mitigates both untargeted and targeted (or backdoor) attacks while demonstrating its effectiveness in non-IID data environments.
-We leave theoretical guarantees of FedCC and experiments when clients have different neural network architectures or tasks as future work.
-Appendix
-Appendix 0.A Dataset and Model Architecture
-We use three benchmark vision datasets: Fashion-MNIST (fMNIST) [32], CIFAR10, and CIFAR100 [14].
-fMNIST consists of 60,000 training and 10,000 testing samples of 28$\times$28 grayscale images and associating labels of 10 classes.
-The labels include T-shirt/top, Trouser, Pullover, Dress, Coat, Sandal, Shirt, Sneaker, Bag, and Ankle boot.
-CIFAR10 consists of 50,000 training and 10,000 testing examples of 32$\times$32 color images and associating labels of 10 classes.
-The classes include airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck.
-CIFAR100 is the same as CIFAR10 except for the categorization.
-It is labeled as 100 classes, each containing 600 images.
-Thus, there are 500 training and 100 testing images per class.
-Each class (i.e., ‘fine’ label) belongs to one of the 20 superclasses.
-Appendix 0.B FedCC: Algorithm
-Algorithm 1 summarizes FedCC’s workflow.
+is an aggregation method that combines local clients’ model parameters based on the Kernel CKA similarities between the global and local models’ PLRs.
+The complete algorithm is provided in Algorithm 1.
+The server first extracts the PLRs from both the global and local models, then computes the RBF Kernel CKA between them.
+We focus on PLRs specifically, as they exhibit the most distinguishable differences between benign and malicious models, as shown in earlier sections.
+The RBF Kernel CKA is chosen for its effectiveness in capturing nuanced similarity differences under non-IID and adversarial settings.
+The resulting similarity values are clustered into two groups using a simple K-means algorithm—though any binary clustering method would suffice.
+Assuming that fewer than half of the clients are adversarial ($n/2$), the larger cluster is designated as the set of ‘candidates.’ This approach does not require manual thresholding; the fixed $K=2$ clustering aligns with the practical assumption that fewer than half of the clients are adversarial, and empirically, this dynamic separation has shown consistent effectiveness across rounds and datasets.
+Given that earlier CNN layers capture global features while later layers encode local ones [16, 26, 24], we apply two types of normalization: within_cka (denoted $b$ in Figure 2), computed within the dominant (larger) cluster, and across_cka (denoted $a$), computed across all clients.
+For the PLR (second-to-last layer), we use within_cka; for all other layers, we use across_cka to compute layer-wise weighted averages.
+This helps preserve the scale of the aggregated parameters, as the weights are normalized to sum to one.
+FedCC is also designed to remain effective against adaptive attacks, including distributed adaptive attacks like DBA.
+Unlike defenses based on outlier detection in parameter space, FedCC evaluates representational similarity using Kernel CKA between PLRs of local and global models.
+This enables the detection of semantic deviations even when malicious updates mimic benign gradients.
+By clustering clients based on CKA similarity and applying soft, layer-specific weighting—rather than hard rejection—FedCC attenuates the influence of suspicious clients without excluding them outright.
+This makes it robust against attacks that aim to blend in with benign behavior, particularly under non-IID conditions where traditional defenses often fail.
+Theoretical Insight.
+While FedCC is primarily supported by empirical evidence, we briefly offer a theoretical perspective on its robustness.
+Let $\mathcal{R}_{i}$ denote the PLR of client $i$, where benign clients’ PLRs are drawn from distribution $\mathcal{D}_{b}$ and malicious clients’ from $\mathcal{D}_{a}$.
+If the inter-client Kernel CKA similarity within $\mathcal{D}_{b}$ is significantly higher than between $\mathcal{D}_{b}$ and $\mathcal{D}_{a}$, then clustering based on PLR similarity can effectively separate benign from malicious clients.
+Prior work has empirically demonstrated that PLRs from benign clients exhibit high intra-distribution similarity across diverse data distributions [31].
+FedCC leverages this separability by softly weighting updates: clients close to the majority cluster receive higher aggregation weights, while those that deviate receive less influence.
+This approach mitigates adversarial updates without relying on hard rejection, and remains effective even when malicious updates mimic benign gradients.
+Therefore, under the mild assumption that benign clients maintain high intra-cluster PLR similarity and form the majority, FedCC is theoretically expected to preserve robustness across a range of attack scenarios.
 Algorithm 1 FedCC
 Input: global_w, n local_w ($w_{1}$, …, $w_{n}$)
-      Output: agg_w, larger_cluster_members
-$\triangleright$ Get PLRs of local models
+      Output: agg_w, larger_cluster_members      $\triangleright$ Get PLRs of local models
 1:for $i<n$ do
@@
 26:return agg_weights, larger_cluster_members
+5 Experiments
+5.1 Dataset and Model Architecture
+We use three benchmark vision datasets: Fashion-MNIST (fMNIST) [33], CIFAR10, and CIFAR100 [15].
+Considering that the end devices are normally incapable of handling heavy computation due to resource or communication constraints, we used lightweight CNN models for experiments.
+Specifically, we use 4-layer, 5-layer CNN for fMNIST and CIFAR10, and LeNet [17] for CIFAR100, as summarized in Table 2.
+Table 2: CNN Architectures for fMNIST, CIFAR-10, and CIFAR-100
+Dataset Layer In Out Ker / Str / Pad Activation
+fMNIST conv2d_1 1 64 $5\times 5$ / 1 / 0 ReLU
+conv2d_2 64 64 $5\times 5$ / 1 / 0 ReLU
+dropout - - 0.25 -
+flatten - 25600 - -
+fc_1 25600 128 - -
+dropout - - 0.5 -
+fc_2 128 10 - -
+CIFAR-10 conv2d_1 3 64 $3\times 3$ / 1 / 0 ReLU
+maxpool2d - - $2\times 2$ / - / 0 -
+conv2d_2 64 64 $3\times 3$ / 1 / 0 ReLU
+maxpool2d - - $2\times 2$ / - / 0 -
+conv2d_3 64 64 $3\times 3$ / 1 / 0 ReLU
+maxpool2d - - $2\times 2$ / - / 0 -
+flatten - 256 - -
+dropout - - 0.5 -
+fc_1 256 128 - -
+fc_2 128 10 - -
+CIFAR-100 conv2d_1 3 6 $5\times 5$ / 1 / 0 ReLU
+maxpool2d - - $2\times 2$ / 2 / 0 -
+conv2d_2 6 16 $5\times 5$ / 1 / 2 ReLU
+maxpool2d - - $2\times 2$ / 2 / 0 -
+flatten - $4$
+fc_1 44944 120 - ReLU
+fc_2 120 84 - ReLU
+fc_3 84 100 - -
+5.2 Non-IID Simulation
+Non-IID is a more feasible assumption considering the diverse and massive nature of data in practice.
+To standardize non-IID settings, we use the Dirichlet distribution, which is widely adopted for modeling client-level label skew [19].
+The concentration parameter $\alpha$ controls the degree of heterogeneity; smaller values lead to clients holding examples from fewer classes.
+We set $\alpha=0.2$ to simulate moderate heterogeneity, following prior works [12, 38, 31].
+This value is commonly used as it balances realism and trainability: smaller values (e.g., $\alpha<0.1$) can induce excessive label imbalance and hinder convergence, while larger values (e.g., $\alpha>0.5$) reduce the heterogeneity to near-IID conditions [5].
+5.3 Experimental Setup and Baselines
+We implement two untargeted attacks from [8] and two targeted attacks from [3] and [34], along with defense baselines including Krum, multi-Krum [4], Coomed [35], Bulyan [11], FLTrust [7], FLARE [31], and FedCC.
+We use ten clients with a 1.0 client participation fraction unless stated otherwise.
+As in [31], we have three adversaries for each untargeted attack and one for the targeted attack.
+In DBA, we randomly select four malicious clients and implant five distributed triggers per batch.
+To compensate for delayed attack effects, we use FedAvg until epoch 30, after which each defense method is applied.
+Note that, for a fair comparison, FLTrust is computed using its previous global model without auxiliary data.
+Benign clients undergo three local epochs, while compromised clients undergo six, with training done using Adam optimizer (learning rate = 0.001).
+5.4 Evaluation Metrics
+We evaluated our defense mechanism using two metrics: backdoor confidence (confidence) and global model test accuracy (accuracy).
+The confidence metric measures misclassification likelihood, ranging from 0 to 1 (lower values are better).
+Accuracy reflects the global model’s overall test performance under defense methods during attacks, ranging from 0 to 100 (higher values are better).
+Both metrics were reported under backdoor attacks, with a focus on accuracy for untargeted attacks, as the attacker’s goal is to reduce overall accuracy.
+6 Results and Discussions
+In this section, we report the performance of our FedCC against two untargeted and two targeted attacks in both IID and non-IID environments.
+We additionally demonstrate the defense effect against a distributed adaptive attack, DBA [34].
+Unlike centralized backdoor attacks, the DBA splits a trigger pattern across multiple adversaries, evading detection by anomaly detection.
+We show the result of CIFAR10, with various defense methods applied.
+6.1 Non-IID Data Environment
+Non-IID is a more feasible assumption considering the diverse and massive nature of data in practice.
+To standardize non-IID settings, we employ Dirichlet distribution that 0 indicates the most heterogeneity, and 100 mimics homogeneity [21].
+A concentration parameter $\alpha$ controls the degree of non-IID; the smaller $\alpha$ is, the more likely the clients hold examples from only one randomly chosen class.
+Since $\alpha$ being 0.2 represents a highly non-IID scenario based on [12, 38], we set it as such, following [38].
+6.1.1 Untargeted Model Poisoning Attacks in Non-IID Setting
+Table 3 provides the global model’s accuracy across three datasets under two untargeted (Fang-Krum and Fang-Med) in a non-IID environment.
+A notable finding is that robust aggregation algorithms often yield lower accuracy than simple averaging (fedAvg).
+This discrepancy arises from the imperfect identification of malicious clients, leading to the erroneous aggregation of their weights along with those of benign clients.
+Under Fang-Krum, Krum, multi-Krum, and Bulyan experience significant accuracy drops (16.51%, 45.88%, and 13.39%, respectively, compared to FedAvg’s 57.14%), reflecting the attack’s design to degrade Krum-based defenses.
+Coomed also struggles, as median weights fail to represent benign clients effectively in non-IID settings.
+FLTrust mitigates Fang-Krum but performs better on CIFAR10 than fMNIST, likely due to fMNIST’s simpler gradients and lower variance, which challenge its ability to distinguish malicious updates.
+FLARE’s lower accuracy stems from its approach of scaling the entire model weights uniformly, which ignores knowledge contributed by individual clients.
+In contrast, FedCC, which selectively averages weights layer-wise, achieves superior performance by boosting or minimizing weights based on their alignment with benign updates.
+FedCC demonstrates the highest accuracy (71.13%, 52.06%, and 14.51% for fMNIST, CIFAR10, and CIFAR100, respectively) and shows significant improvement in CIFAR100 experiments.
+Turning to the Fang-Med attack, FedAvg experiences a significant drop in accuracy due to the attack’s creation of outliers by deviating from median values.
+Consequently, Coomed and Bulyan perform relatively well since their coordinate-wise median and trimmed mean methods disregard outliers effectively.
+FLARE continues to perform poorly for a similar reason mentioned earlier.
+Between the Krum-based methods, multi-Krum and Coomed are more robust than Krum, as they aggregate multiple local models, providing better resilience to outliers.
+Similarly, Bulyan’s averaging of multiple clients enhances robustness but falls short of Coomed, particularly in non-IID environments.
+Finally, FedCC achieves the highest accuracy (72.76%, 47.85%, and 16.12% for each dataset) without sharing raw data, showcasing its exceptional performance.
+This success is attributed to two factors: (1) highly distinguishable information within PLRs and (2) higher CKA similarity between benign clients than between benign and malicious clients.
+Table 3: Test Accuracy under untargeted attacks in Non-IID setting.
+Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
+Fang -Krum non-IID fM 57.14 16.51 45.88 57.12 13.39 60.8 49.54 71.13
+C10 33.69 15.38 20.5 35.7 19.23 41.87 17.03 52.06
+C100 2.27 1 4.95 7.85 0.98 11.04 7.46 14.51
+Fang -Med non-IID fM 16.32 49.33 66.84 68.9 64.12 18.96 52.25 72.76
+C10 10.02 25.06 45.44 40.23 32.47 10 14.59 47.85
+C100 1 6.24 14.52 10.27 6.91 1.09 1 16.12
+6.1.2 Targeted Backdoor Attacks in Non-IID Setting
+Table 4 summarizes test accuracy under targeted attacks in a non-IID setting.
+To further illustrate the impact, Figure 4 provides a visual representation of backdoor confidence.
+It is evident that Krum, which selects a single client’s weights as the global model, shows reduced robustness compared to methods that aggregate weights from multiple clients.
+In contrast, Coomed, multi-Krum, and Bulyan exhibit higher test accuracy.
+FLARE reduces main task accuracy for reasons similar to those observed in previous experiments.
+FedCC achieves the highest test accuracy.
+Since targeted attacks aim to misclassify a specific target class while maintaining the main task accuracy, the best performance of FedCC is attributed to leveraging knowledge from earlier stages of training.
+This result highlights the superiority of using sophisticated weighted knowledge and filtering malicious clients via CKA similarity, which outperforms first-order statistical methods like mean or median.
+A similar trend is observed with the DBA attack, except for multi-Krum; the lowest accuracy occurs because all four malicious clients, each with a distributed trigger, are treated as outliers and excluded from aggregation.
+In contrast, FedCC preserves high main task accuracy (52.28%) even under DBA, outperforming all baselines.
+This result highlights its robustness against adaptive strategies that distribute triggers across multiple compromised clients—a setting specifically designed to evade anomaly-based filtering.
+Table 4: Test Accuracy under targeted attacks in Non-IID setting.
+Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
+Target non-IID fM 75.65 45.27 65.97 71.70 57.96 61.82 64.31 75.66
+C10 36.16 14.98 30.72 48.97 40.11 44.06 10.18 51.56
+C100 4.46 6.18 6.90 12.04 11.16 12.95 1.14 15.26
+DBA C10 38.56 24.94 7.09 44.45 34.19 51.49 38.73 52.28
+Notably, Figure 4 demonstrates that FedCC significantly reduces backdoor confidence.
+Note that since the backdoor confidence of DBA fluctuates a lot due to its distributed nature, we omit it for brevity.
+Unlike other methods, such as FLTrust or multi-Krum, which exhibit fluctuating confidence values, FedCC maintains consistently low confidence, emphasizing its resilience and independence from client-specific data distributions.
+This advantage stems from the effective utilization of CKA, enabling similarity calculations even when models are trained on different datasets with varying distributions.
+While other baselines, like Bulyan or FLTrust, also reduce confidence, FedCC not only mitigates the backdoor attack but also sustains—or improves—accuracy by precisely identifying and down-weighting malicious clients.
+Figure 4: Confidence of Backdoor Task for targeted attacks in Non-IID settings.
+Summary of Non-IID Robustness.
+Across all attack types, FedCC consistently demonstrates superior performance in non-IID environments—outperforming prior defenses in both main task accuracy and backdoor confidence.
+This robustness stems from the use of Kernel CKA over penultimate layer representations, which captures distributional structure even under client heterogeneity.
+Unlike methods based on geometric or statistical outlier filtering, FedCC’s representation-based, layer-aware weighting enables it to preserve benign signals while attenuating subtle or distributed malicious updates, making it particularly effective under realistic non-IID conditions.
+6.2 IID Data Environment
+In an IID setting, we evenly divide the training dataset among all clients so that each client’s data distribution is identical and of the same size.
+6.2.1 Untargeted Model Poisoning Attacks in IID Setting
+Table 5 shows the test accuracy of the global model trained on three datasets under untargeted attacks in IID settings.
+While trends in these experiments are similar to those in the non-IID setting, the accuracy values are slightly higher due to the more consistent data distribution across and updates from clients in the IID case.
+A notable deviation from the non-IID experiments is when Coomed is applied: the accuracy increases from 75.55% (FedAvg) to 87.62% (Coomed).
+This indicates that the coordinate-wise median values closely align with the global model’s weights with minimal deviation and represent benign models well, as the adversary’s impact is limited due to the majority of benign clients.
+FedCC achieves the highest accuracy across all methods (89.57%, 69.84%, and 18.47% for each dataset) due to its effective CKA similarity measurement and sophisticated layer-wise aggregation strategy.
+Under the Fang-Med attack, the accuracy of FedAvg experiences a significant drop (89.89% to 20.86%) due to the peculiarity of outliers when the clients have IID data.
+Similar to the result of Fang-Krum, Krum, Coomed, and Bulyan demonstrate similar accuracy values, indicating that coordinate-wise median values indeed represent benign clients’ parameters.
+FLTrust, however, is less effective in mitigating the Fang-Med attack than Fang-Krum, even in the IID setting; it then means that the performance difference is not due to data distribution but rather its inherent vulnerability.
+In Fang-Krum, malicious updates are deliberately far from the global model, allowing FLTrust to identify discrepancies easily.
+In contrast, the malicious updates in Fang-Med are subtler and only slightly deviate from the global model or median, making it harder for FLTrust to distinguish them from benign updates.
+Meanwhile, FedCC consistently achieves the highest accuracy (89.66%, 70.52%, and 17.83% for each dataset), demonstrating its capability to mitigate attacks involving more nuanced changes to model updates without extreme outliers.
+Table 5: Test Accuracy under untargeted attacks in IID setting.
+Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
+Fang -Krum IID fM 75.55 31.66 87.78 87.62 50.30 89.53 79.16 89.57
+C10 49.67 40.86 63.42 57.40 12.67 68.25 25.77 69.84
+C100 13.72 1.04 7.64 6.17 1.59 17.14 7.49 18.47
+Fang -Med IID fM 20.86 85.33 89.53 86.70 87.45 21.36 71.08 89.66
+C10 9.51 54.28 69.68 59.20 57.69 9.92 49.82 70.52
+C100 0.87 12.27 16.52 14.43 12.56 1.16 5.83 17.83
+6.2.2 Targeted Backdoor Attacks in IID Setting
+Table 6 and Figure 5 present the test accuracy and backdoor confidence under targeted backdoor attacks in IID settings.
+Similar to the untargeted attack scenario, the general test accuracy in IID settings is higher than in non-IID settings.
+Reducing backdoor confidence is as critical as maintaining main task accuracy, as attackers aim to stealthily embed backdoor tasks.
+A key observation is FedCC’s remarkable ability to reduce backdoor confidence to near zero while preserving high main task accuracy.
+In contrast, fluctuating backdoor confidence observed in Multi-Krum and Flare is largely due to the distributed trigger, which confuses their defense mechanisms and hampers their ability to filter outliers effectively.
+Other methods, such as Multi-Krum, Flare, and FLTrust, show significant variability in backdoor confidence, further highlighting FedCC’s superiority.
+Notably, FedCC achieved the highest main task accuracy across both targeted backdoor attacks and DBA, consistently neutralizing the backdoor task.
+These results underscore FedCC’s ability to mitigate targeted backdoor attacks through precise and efficient integration of prior knowledge to preserve main task performance while zeroing out the attack confidence.
+Table 6: Test Accuracy under targeted attacks in IID setting.
+Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
+Target IID fM 88.27 86.63 87.03 89.41 89.45 89.59 75.29 90.01
+C10 64.68 57.69 71.19 69.85 68.76 68.61 11.09 71.64
+C100 13.83 5.72 17.57 13.24 15.11 17.08 1.03 18.61
+DBA C10 10.00 35.58 51.40 10.00 16.16 56.69 10.00 58.04
+Figure 5: Confidence of Backdoor Task for targeted attacks in IID settings.
+6.3 Performance in Varying FL Settings
+We additionally measured the effectiveness of FedCC in various non-IID settings, including different numbers of malicious clients, fractions of participation, and numbers of local epochs.
+Since Coomed was as effective as ours, and FLARE also used PLRs, we compared three methods to FedCC, including FedAvg.
+6.3.1 Impact of the Number of Malicious Clients
+We assessed the impact of different numbers of malicious clients under untargeted attacks, specifically Untargeted-Krum and Untargeted-Med.
+In this set of experiments, we kept the total number of clients and participation fraction fixed at 10 and 1, respectively.
+To adhere to the assumption that the number of attackers is less than half of the participating clients, we investigated the test accuracy considering scenarios involving up to four malicious clients.
+Figure 6 illustrates the accuracy in the given experimental settings, and it is evident that FedCC outperforms the other defense methods.
+We observe a general trend where the accuracy of defense methods decreases as the number of attackers increases.
+Notably, Coomed experiences the most significant drop in accuracy when transitioning from one malicious client to four malicious clients.
+This can be attributed to the fact that as the participation of malicious clients increases, there is a higher likelihood of the median values being influenced by their malicious contributions.
+In contrast, FedCC does not rely solely on geometric measures such as angles or distances but instead leverages hidden correlations between networks to identify and filter out malicious clients.
+As a result, FedCC demonstrates superior accuracy in filtering out malicious clients, regardless of the number of attackers involved.
+It is important to note that the accuracy degradation observed as the number of malicious clients increases primarily due to aggregating fewer clients rather than the malicious clients themselves being selected.
+These findings underscore the robustness and effectiveness of FedCC in defending against untargeted attacks, as it consistently outperforms other defenses across varying numbers of malicious clients.
+Figure 6: Test Accuracy with different numbers of malicious clients
+6.3.2 Impact of Fraction
+We examine the impact of the participating fraction of clients where the number of total clients is fixed at 100.
+The fractions are 0.1, 0.3, and 0.5, such that the numbers of clients being aggregated are 10, 30, and 50, respectively.
+The proportion of malicious clients is fixed at 0.3, such that the malicious clients are 3, 9, and 15, respectively.
+Table 7 summarizes the accuracy.
+We can observe that the accuracy tends to increase as the participation fraction is greater under untargeted-Krum attacks.
+It implies that the more client participation in a non-IID setting, the more accurate the global model is if the attacks are mitigated.
+Under untargeted-Med attacks, the accuracy is the highest with FedCC and second highest with Coomed.
+Throughout the experiment, FLARE prevented the global model from convergence, and the test accuracy fluctuated.
+The fluctuation is presumably due to the weight scaling while randomly selecting clients.
+Different from previous experiments, the server chooses a fraction of the client every round, meaning not-yet-trained models can be selected.
+In this environment, weight scaling leveraged by FLARE constantly improperly scales the local weights, causing divergence and fluctuation.
+These experimental results indicate that selectively choosing clients based on CKA similarity is more reliable than taking the median value of each coordinate in a non-IID setting.
+We also observed that as the participation fraction grows, the accuracy grows.
+This is due to the rising number of aggregated local models; the more models are aggregated, the more general global model is generated.
+Table 7: Test Accuracy under untargeted attacks with different fractions of clients being selected.
+Untargeted-Krum Untargeted-Med
+Frac Data FedAvg Med FLARE FedCC FedAvg Med FLARE FedCC
+0.1 fM 55.31 49.83 34.02 64.83 16.57 66.41 52.24 69.52
+C10 10.06 22.55 14.50 20.49 10.00 15.33 10.00 29.81
+0.3 fM 64.22 57.52 10.00 73.55 16.26 58.07 10.00 61.12
+C10 24.24 12.59 10.00 27.81 10.98 22.61 10.00 38.27
+0.5 fM 62.37 58.20 10.00 76.41 18.36 62.49 10.00 69.05
+C10 23.99 17.28 10.06 34.32 9.87 27.83 10.00 37.27
+7 Limitation
+While FedCC demonstrates strong performance against various poisoning attacks in both IID and non-IID settings, several limitations remain.
+First, our experiments focus on lightweight CNN architectures (e.g., LeNet and custom CNNs) and relatively small-scale vision datasets such as fMNIST, CIFAR10, and CIFAR100.
+While these choices reflect practical FL deployments on resource-constrained devices, the generalizability of FedCC to larger-scale datasets (e.g., ImageNet) or more complex architectures (e.g., ResNet, ViTs) remains untested.
+Additionally, FedCC currently assumes homogeneous model architectures across clients.
+Extending it to heterogeneous or multi-task client settings is an important direction for future work.
+Second, while FedCC does not rely on explicit access to raw data and has shown strong empirical robustness, including against adaptive attacks like DBA, it lacks formal guarantees under broader adversarial conditions.
+Theoretical analysis of its robustness bounds, especially under targeted mimicry of benign PLR distributions, remains open.
+Our theoretical insight provides a preliminary foundation but invites further exploration.
+Third, while Kernel CKA offers strong representational comparison performance, it incurs computational overhead during aggregation.
+Although this cost is manageable in our experimental setting, optimizing its computation or exploring more efficient alternatives would be necessary for scalability in large-scale deployments.
+8 Conclusion and Future Work
+FL has emerged in response to growing concerns about data privacy in collaborative machine learning.
+However, the distributed nature of FL introduces vulnerabilities to model poisoning attacks, especially under non-IID data settings.
+While many existing defenses are effective under specific threat models, they often neglect the challenges introduced by client heterogeneity.
+In this paper, we proposed FedCC, a robust aggregation method that leverages Kernel CKA to measure representational similarity in the penultimate layer of client models.
+Through extensive experiments across multiple datasets and attack types, we demonstrated that FedCC effectively mitigates both untargeted and targeted (backdoor) attacks, even under severe non-IID conditions.
+For future work, we plan to provide theoretical guarantees for FedCC’s robustness and extend it to scenarios with heterogeneous model architectures.
+We also aim to evaluate its generalizability on larger and more complex datasets (e.g., ImageNet) and explore its effectiveness against stealthier or adaptive attack strategies.
+Lastly, we will investigate optimization techniques to reduce the computational cost of Kernel CKA, enabling scalable deployment in large-scale federated systems.